A Complete Bibliography of Publications in WIREs Computational Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 April 2025 Version 1.10

Title word cross-reference

g [434]. h [524]. K [375, 403]. L_{1} [406]. L_{q} [300]. N [610, 455]. p [455]. U [176].

-likelihood [524]. -mixture [610]. -omic [333]. -regularization [406]. -sample [403]. -small- [455]. -statistics [176].

10-year [236]. 19 [602].

21st [44].

absolute [240]. abundance [577]. accelerated [644]. Accelerating [426]. Acceleration [609]. accessibility [528]. accessible [632]. ACM [41]. activation [243]. Actuarial [117]. acyclic [467]. Adaptations [660]. adapted [493]. Adaptive [311, 16, 604]. additional [390, 410]. additive [591]. Advance [515]. Advances [539, 613, 400, 494, 466, 639, 587, 394].

Advancing [622]. Adversarial [537, 487, 615]. Africa [19]. Agent [165, 348]. Agent-based [165, 348]. Aggregated [438]. Aggregating [509]. aggregation [108]. Akaike [446]. Alarm [243]. Algorithm [74, 591, 609, 493]. Algorithms [16, 41, 453, 647, 242, 426, 245, 248, 592]. Alternating [281]. alternative [347]. Alzheimer [599]. Analysis [479, 35, 463, 96, 198, 258, 448, 537, 159, 458, 239, 174, 251, 625, 533, 374, 455, 229, 483, 494, 616, 197, 114, 213, 648, 514, 585, 146, 575, 521, 532, 260, 377, 226, 621, 5, 427, 421, 207, 552, 525, 507, 317, 389, 428, 217, 534, 223, 321, 604, 645, 628, 327, 333, 404]. Analytical [399]. analytics [606, 556]. Analyzing [658, 325, 481, 212, 278]. Andrews [30, 167]. angle [294]. Animal [495]. animation [43]. Anomaly [344, 243]. Antithetic [15]. Applet [667]. application [446, 513, 274, 602]. Applications [661, 662, 377, 657, 155, 544, 444, 543, 549, 509, 615, 202, 562, 453, 150, 619, 540, 176]. Applied [17, 347]. Applying [305]. approach [100, 617]. Approaches [650, 398, 598, 516, 599, 612, 484, 629, 518, 631, 492, 379]. approximate [482]. Approximation [307, 60]. AQ [74]. architecture [12]. architectures [82]. Archive [221]. area [608, 436]. area-level [608]. areas [192]. arm [625]. arm-based [625]. art [558]. Artificial [199]. artistic [197]. aspects [557, 488]. assessment [399, 451]. associated [456]. association [251, 411, 375, 589]. asymmetrical [250]. asymptotic [620]. asymptotics [244]. attribution [579]. authorship [579]. autocorrelation [343]. Autocovariance [24, 416]. automata [298]. Automated [317]. Autoregressive [163, 513, 369, 289, 493, 654]. Average [654, 369]. Averaged [68]. axioms [424].

B [75]. B-splines [75]. backfitting [591]. Background [446, 202]. balance [282]. balancing [311]. bands [385]. bandwidth [475]. Based [658, 660, 656, 591, 165, 348, 625, 369, 506, 643, 484, 629, 414, 563, 241, 411, 20, 257, 275, 304, 450, 217, 319, 325, 582, 393, 528, 447]. bases [262]. Bayes [443]. Bayesian [191, 417, 398, 353, 506, 100, 349, 341, 482, 219, 234, 235, 121, 40, 530, 518, 436, 310, 293, 521, 573, 574, 522, 377, 202, 408, 233, 434, 452, 428, 568, 617, 357, 520, 619, 327, 468]. BDeu [520]. be [318]. Beamlets [63]. behavior [325]. between [530, 19]. beyond [534]. Bias [301]. biasedness [588]. Big [606, 455, 330, 305, 597]. Bigram [78]. Bill [265]. Bin [220]. binary [625]. bio [541]. bio-inspired [541]. Bioavailability [312]. bioequivalence [312]. Biology [666, 401, 230]. Biomarker [627]. biomedical [165]. Biometric [267]. Biplots [328, 366, 222]. birth [412]. birth-death [412]. bivariate [471]. blind [555]. Boosting [58, 607, 295]. Bootstrap [180, 475, 206]. Bootstraps [276]. boundary [462]. Bounds [109]. Box [148]. Breakdown [38]. Breaking [666]. bridge [42]. brief [505, 270, 526]. Brownian [42, 145]. Brushing [126]. BSiZer [105]. bubbles [635]. building [223].

calculations [172]. calibrated [607]. Calibration [361, 630]. can [592].

Cancer [666, 289]. capture [255]. capture-recapture [255]. Carlo [637, 271, 586, 235, 623, 190, 320, 307, 481, 527]. case [132, 319]. catch [70]. Cauchy [460]. causal [573, 576]. cell [556]. cellular [298]. Censored [71, 580]. Centroids [33]. century [44]. chain [271, 481]. Challenges [660, 257, 534, 267, 615, 634, 430, 487]. change [349, 451, 282, 614]. change-point [614]. changes [430]. character [526]. choices [628]. Cholesky [565, 31]. chromatin [528]. Ciphers [157]. Circular [99]. Class [70, 641, 282]. class-balance [282]. classes [374]. classical [403, 267]. Classification [440, 501, 433, 484, 365, 624]. classifiers [89]. cleaning [147]. climate [218, 451, 430, 493]. clinical [172, 55]. CLOUD [136, 410]. Cluster [585, 552, 217]. Cluster-scaled [552]. Clustering [660, 486, 272, 396, 571, 353, 465, 643, 433, 618, 569, 522, 603, 257, 550, 447]. Clusters [192, 553]. co [353]. co-clustering [353]. Coding [34]. coefficients [369, 125]. collaborative [194]. Collected [41]. Collections [667]. Color [142]. colorectal [289]. Combining [608, 47, 184]. Commander [237]. commonalities [483]. communication [617]. Community [544, 324, 639, 414, 394]. community-based [414]. comparison [625, 612, 186]. comparisons [408]. Competing [85, 525]. competitions [350]. completion [457]. complex [544, 278, 359, 325, 545]. Complexity [666]. component [96, 198, 258, 239, 427, 552]. components [281, 462]. comprehensive [626, 615, 221]. Computation [49, 456, 439, 614, 7, 482]. Computational [80, 412, 218, 314, 650, 306, 230, 137, 20, 538, 557, 666, 488, 273, 1, 3, 457, 315, 346, 150, 540, 215]. Computations [159, 274, 35]. Computer [107, 135, 64, 195, 630, 487]. computing [505, 318, 629, 136, 207, 410, 299, 237]. concave [439]. concepts [219, 35]. concurrent [383]. Conditional [262, 499, 477, 289]. confidence [385, 154]. confidentiality [175]. Conjugate [46]. connect [227]. connections [206]. considerations [589]. Constrained [391, 448, 601]. Constructing [420]. consumer [279, 410]. Content [357]. Context [357]. contingency [262, 461]. Continuous [43, 500, 601, 263, 477, 520]. contrast [625]. contrast-based [625]. control [593]. converge [315]. Convergence [647, 522]. Conway [531]. cooperation [617]. coordinate [144]. coordinates [229, 123]. Coping [141]. Copulae [551]. Correction [637, 203]. correlates [599]. correlation [314, 442]. correlations [277]. Correspondence [114, 458]. count [386, 410, 531, 498, 540]. counterparts [227]. counts [568, 577]. courses [227]. Covariance [266, 159, 496, 422, 53, 415, 401, 473, 508, 467]. covariances [308]. covariate [612, 282]. covariates [311]. Cover [535, 503, 363, 70]. COVID [602]. COVID-19 [602]. Cox [212]. crave [604]. crew [410]. Crisis [653]. criteria [596]. criterion [446, 202]. Critical [541]. cross [432, 187]. cross-disciplinary [187]. Crossover [115]. cultures [636, 316]. Current [487, 288, 494]. curve [475, 391]. curves [479, 167]. CXXR [259]. cybersecurity [487]. cycle [410]. Cytometry [660, 521].

Data [155, 632, 658, 501, 10, 455, 132, 514, 175, 576, 589, 379, 258, 252, 268, 51, 479, 239, 396, 134, 583, 287, 580, 373, 533, 203, 385, 544, 374, 309, 218, 556, 519, 486, 71, 318, 328, 366, 213, 444, 612, 622, 386, 291, 484, 607, 414, 500, 594, 575, 436, 99, 558, 521, 532, 289, 641, 451, 330, 377, 5, 278, 352, 421, 525, 531, 317, 389, 406, 450, 420, 362, 577, 133, 147, 534, 308, 223, 321, 453, 602, 627, 645, 624, 493, 150, 332, 619, 545, 463, 327, 550, 333, 393, 447, 305, 597]. DataDesk [223]. datasets [191, 4, 141, 559]. DC [19]. death [412]. debiasing [604]. decades [263]. Decision [280, 359]. decomposition [448, 472, 222, 302]. Deep [636, 488, 230, 645]. defense [424]. defines [318]. Delaunay [109]. Delta [620]. denoising [455]. densities [574, 247]. density [423, 98, 482, 13, 144, 439, 453]. dependence [52, 562]. dependency [283]. derivation [446, 202]. Derived [656, 386]. descent [128]. design [265, 110, 29, 382, 142]. Designs [665, 564, 185, 139, 16]. Detecting [553]. Detection [8, 635, 203, 544, 324, 466, 344, 634, 243, 526, 614, 394]. determinant [53, 415]. determining [599]. Development [661, 312]. Developments [395, 659, 551, 278, 487, 254]. Diagnostic [249, 28, 139, 522]. differences [625, 530]. Differential [533, 507, 581, 367, 273, 577]. digraphs [70]. dimension [472, 329, 515, 390, 450, 340]. dimensional [502, 287, 519, 473, 558, 421, 406, 179, 467, 308, 453, 582, 440]. dimensionality [141]. dimensions [260]. directed [467]. directions [355]. disciplinary [187]. disclosure [10]. Discounting [134]. discovering [256]. discovery [151, 243]. Discrete [657, 641, 277, 525, 520]. Discriminant [226, 648, 260, 421]. discrimination [334]. Diseases [594]. disparities [619]. dispersed [531]. display [112, 223]. Distance [334, 447, 360]. Distance-based [447]. Distance-weighted [334]. DISTATIS [198]. Distributed [82]. Distribution [25, 657, 620, 386]. Distribution-free [25]. distributions [416, 471, 386, 291, 161, 601, 263, 290, 557, 602]. DNA [268]. do [315]. Document [571, 288]. domain [433, 379]. Dominated [242]. Dot [250]. down [400]. dreams [292]. Drug [661, 312]. Durban [19]. dynamic [213, 558, 344, 562]. dynamical [241].

Early [653]. easy [227]. Echelon [575]. econometrics [338]. economics [398]. Ecosystem [651]. edge [629]. editing [203]. Editor [284, 381]. Editorial [511]. Education [667]. educational [543, 549]. effect [462, 254]. effects [200, 301, 238]. efficient [564, 203, 527]. elastic [479]. Electroencephalogram [269]. Electroencephalogram-sleep [269]. electronic [534]. elimination [156]. ellipsoid [9]. else [237]. emission [599, 213]. Empirical [443, 324, 612, 587, 641, 391]. Endpoints [661]. Energy [360, 513]. engineering [264]. ensemble [295, 248]. ensembles [283]. Envelope [470]. environment [621]. epidemiological [594]. epidemiology [322]. equating [549]. equation [310]. equations [581, 533, 274, 367, 273]. era [597]. Erratum [178, 118]. Error [593, 559, 627]. errors [203]. estimate [439]. Estimating [214, 430]. Estimation [422, 561, 401, 308, 240, 591, 475, 496, 398, 423, 323, 343, 98, 69, 608, 482, 219, 234, 235, 607, 301, 473, 569, 13, 538, 613, 275, 467, 319, 570, 453, 619, 409]. evaluation [350, 324, 186, 195, 72]. event [95, 525, 359, 212]. Evolution [318, 526]. Evolutionary [7]. exact [456]. Examining [291]. Examples [252]. exchangeability [639]. expectation [278, 493]. expectation-maximization [278]. Experimental [3, 265, 564, 110, 382]. Experimentation [665, 256]. experiments [115, 509, 604]. experts [509]. explainable [598]. exploration [499, 223]. Exploratory [5, 616]. explorer [194]. exponential [471]. exponential-geometric [471]. exponentiated [471]. expression [528]. extensible [259]. extensions [415]. extraction [351]. extremes [539, 638].

face [267, 451]. Factor [448, 665, 258, 253, 616]. Factorial [185, 29]. factorization [31]. factorizations [506]. families [263, 277]. farm [410]. faster [276]. fatality [377]. feature [444]. features [268, 440]. fiducial [296]. Field [655, 502]. Filtered [13]. Filtering [484, 428]. Filtering-based [484]. filters [234, 235, 86, 404]. final [628]. finance [244, 314, 162]. Financial [290, 22, 184, 635, 562, 389, 321]. finding [349]. finite [618]. fit [170, 461]. fitting [255, 342]. Fixed [200]. Flow [660, 502]. fMRI [455]. forecasting [502, 513]. forecasts [186]. forensic [21]. formal [605]. formulations [448]. Forward [57]. foundations [544]. Fourier [369]. Fractal [145, 208]. Fractional [29, 145, 347]. frailty [494]. Frameworks [331]. free [25, 639, 331, 147]. frequency [433, 321]. frequentist [398, 530]. Function [567, 416, 409]. Functional [597, 238, 645, 513, 385, 626, 484, 532, 383, 613, 450, 624, 327]. functions [88, 508, 333]. Fundamentals [615]. fusion [489]. future [267, 634, 365]. Fuzzy [84].

Game [207]. Game-theoretic [207]. Gauss [224]. Gaussian [373, 234, 156, 467, 570]. gene [621, 628, 256, 528]. gene-environment [621]. general [219]. generalization [390]. Generalized [296, 650, 171, 524, 607, 409]. generating [242, 263]. generation [639]. generative [615]. Generators [659, 395]. genetic [468]. Genome [528, 589]. Genome-wide [528, 589]. genomic [598]. genomics [492]. GenStat [32]. Geographic [23, 118]. geometric [471, 500]. Geometry [124, 208, 229]. Geospatial [81]. geostatistical [486]. ggplot2 [149]. Gibbs [283, 547]. Gini [416]. given [262, 467]. Global [168, 229]. goes [265]. Goodness [461, 170]. Goodness-of-fit [461]. gradient [46]. grammar [122]. Grand [30, 261]. graph [467]. Graphical [306, 143, 252, 490, 570]. Graphics [27, 64, 122]. graphs [335, 231]. gravitational [538]. grid [82]. grid-type [82]. Gromping [424]. group [518]. groups [530]. guide [351, 628]. GWAS [589].

HALT [644]. Hanging [250]. Harmonic [146]. Harnessing [414].

Hastings [647]. Hate [634]. health [279, 436, 534, 619]. Heavy [247, 416, 290]. heavy-tail [416]. Heavy-tailed [247, 290]. Hidden [662, 323]. Hierarchical [396, 322, 524, 289, 249]. High [473, 467, 502, 287, 519, 141, 558, 260, 421, 406, 308, 321, 453, 582, 440]. High-dimensional [473, 467, 502, 287, 519, 558, 421, 308, 453, 582]. high-frequency [321]. Higher [244, 345]. Higher-order [244, 345]. Highly [644]. Histogram [54, 396, 68]. historical [613]. History [210, 270, 293, 526]. homogeneity [291]. homogeneous [123]. homology [512]. horseshoe [452]. Hubert [250]. Hunter [265]. hybrid [468]. hyper [434]. hyper- [434]. hyperspectral [466]. hypotheses [492]. Hypothesis [405, 593].

ideas [606, 265]. identifiability [323]. Identification [268, 95, 641, 555]. identifying [250]. II [418, 234, 549, 545]. III [235]. illustrative [580]. Image [158, 535, 261, 14, 503, 363, 213]. imaging [599, 466, 327]. implementation [543]. Implicit [188]. Importance [56, 371, 336, 424]. important [320]. Improving [547]. Imputation [246, 493, 393]. imputation-based [393]. incompatible [283]. incomplete [607]. independence [477]. Independent [427]. Individualized [489]. individuals [495]. Inference [482, 462, 489, 417, 581, 586, 100, 561, 213, 438, 296, 262, 573, 331, 613, 233, 104, 304, 576]. inflated [540, 545]. informatics [279]. Information [376, 378, 380, 384, 388, 392, 397, 402, 407, 413, 419, 425, 431, 437, 441, 445, 449, 454, 459, 464, 469, 474, 480, 485, 491, 497, 504, 510, 517, 523, 529, 536, 542, 548, 554, 560, 566, 572, 578, 584, 590, 595, 600, 611, 633, 640, 642, 646, 649, 652, 664, 93, 59, 596, 155, 216, 225, 358, 364, 368, 370, 372, 446, 390, 262, 382, 202, 47, 23, 118]. INLA [435]. inspired [516, 541]. Institute [17]. Instructional [667]. Insurance [451]. Integer [48]. Integer-valued [48]. Integrating [451]. integration [576]. Integrative [550, 333, 492]. intelligence [199]. interaction [621]. interactive [223]. interface [318, 346]. interfaces [143]. interpolation [641]. interpretable [598]. interpretation [446]. interpreter [259]. interval [239, 362]. interval-valued [362]. intervals [563]. Intervention [653]. Introducing [1]. Introduction [228, 512]. introductory [227, 304]. intuitive [227]. Inverse [417, 450, 345, 607]. investigation [606]. irregular [192]. Issue [216, 225, 358, 364, 368, 370, 372, 376, 378, 380, 384, 388, 392, 397, 402, 407, 413, 419, 425, 431, 437, 441, 445, 449, 454, 459, 464, 469, 474, 480, 485, 491, 497, 504, 510, 517, 523, 529, 536, 542, 548, 554, 560, 566, 572, 578, 584, 590, 595, 600, 611, 633, 640, 642, 646, 649, 652, 664, 503, 363]. issues [589]. Item [543, 549].

jackknife [206]. James [516]. JMP [151]. job [237]. Joint [570]. Journal [215, 18, 49, 306]. journey [626]. joy [222].

Kalman [234, 428]. Kaplan [391]. Kernel [411, 423, 230, 13].

Kernel-based [411]. Kinetic [213]. knots [119]. Kolmogorov [86, 404].

L2E [6]. landmarks [479]. language [87]. large [191, 455, 324, 335, 461]. large- [455]. large-scale [324]. Lasso [592]. Latent [448, 650, 61, 436]. lattice [575]. lattices [192]. law [635]. Learning [631, 352, 282, 653, 152, 636, 489, 74, 287, 598, 400, 228, 182, 230, 488, 304, 520, 150, 487, 354]. Least [294, 61, 281, 73, 567, 342, 325, 245]. Letter [284]. Level [429, 608]. Leveraging [330]. LIBRA [103]. library [103]. life [644]. lifetime [580]. Likelihood [104, 524, 587, 439, 409, 391]. limitation [10]. limitations [625, 605]. limits [356, 602]. Linear [650, 209, 362, 240, 286, 519, 524, 518, 249, 613, 171, 462, 409, 205]. linguistics [346]. link [409]. Linkage [656, 106, 559, 313]. linked [583, 559]. Lisp [116]. Lisp-Stat [116]. literature [612, 627]. load [410]. location [442]. Loess [111]. Log [205, 635, 439]. log-concave [439]. Log-linear [205]. log-periodic [635]. logistic [460]. logit [251]. Long [52]. Long-range [52]. longitudinal [287]. LPPLS [635].

Machine [152, 653, 150, 598, 400, 178, 153, 487]. machines [36, 154, 420]. magnetic [327]. magnitude [276]. majority [354]. making [359]. Management [653, 410, 562]. manifold [228]. Manifolds [229]. many [227, 389]. mapping [442]. maps [349, 594]. margins [295]. mark [236]. Markdown [339]. markets [314]. Markov [271, 323, 262, 481]. masking [10]. Mass [660]. Massive [4, 373, 141]. Matching [313]. Mathematical [17, 325]. mathematics [3, 184, 315]. MATLAB [103]. MATLAB(R) [137]. matrices [565, 455]. Matrix [457, 448, 274, 401, 473, 194, 78]. maximization [278, 493]. maximum [439, 409]. Maxwell [531]. MCMC [426]. mean [385, 607]. meaningful [382]. measurement [543, 549, 627]. measurements [628]. measures [183, 411, 59]. mechanisms [291, 639]. media [634]. median [341]. medical [627]. Meier [391]. melanoma [466]. Merging [636]. meta [625]. meta-analysis [625]. metamodeling [325]. method [620, 206, 483, 320, 641, 138, 177, 46, 6, 117, 224]. Methodology [214, 66, 166]. Methods [658, 263, 656, 83, 252, 272, 443, 181, 418, 287, 189, 25, 227, 470, 412, 203, 369, 218, 400, 120, 102, 482, 444, 433, 67, 622, 390, 429, 190, 108, 230, 603, 73, 509, 377, 621, 278, 555, 195, 450, 211, 577, 359, 245, 45, 619, 256, 527, 550, 393]. metric [198]. Metropolis [647]. microarray [268]. microbiome [577]. microdata [622]. mind [617]. Mine [632]. minimization [567]. Minimum [53, 415, 9, 561, 204]. mining [155, 51, 351, 150, 379, 579]. Minitab [129]. missing [203, 309, 612, 291, 420, 393]. mixed [238, 462, 447]. mixture [618, 521, 610, 602]. mixture-of-normal [602]. Modal [423]. Model [660, 499, 217, 147, 632, 341, 524, 310, 289, 574, 432, 257, 408, 233, 635, 630, 325, 570, 223, 212, 294, 596]. Model-Based [660, 217, 257]. Model-free [147]. Modeling [638, 264, 214, 165, 435, 191, 184, 367, 539, 322, 436, 290, 298, 166, 562, 568, 617, 540, 545, 468, 205]. Models [654, 655, 662, 650, 187, 240, 591, 371, 348, 496, 255, 251, 286, 513, 323, 373, 253, 506, 608, 494, 482, 336, 200, 618, 607, 490, 563, 500, 518, 241, 521, 238, 249, 610, 383, 613, 171, 462, 495, 531, 214, 467, 57, 546, 498, 248, 256, 540, 545, 409, 327]. Modern [403, 631, 656, 332, 527, 585, 597]. Mollified [657]. Moment [275, 450]. moment-based [450]. Mondrian [113]. monotone [375]. Monte [637, 271, 586, 235, 623, 190, 320, 307, 481, 527]. Mosaic [201]. motion [145]. movement [495]. Moving [654, 369]. Muller [148]. multi [612, 550]. multi-omics [612, 550]. multiblock [258]. Multicollinearity [90]. multidimensional [198]. multifractal [208]. multilevel [490]. Multinomial [656, 251]. Multinomial-Derived [656]. Multiple [258, 458, 490, 492, 408, 495]. multiplexing [356]. Multiscalar [468]. multiscale [506]. multitable [198, 258]. Multivariate [189, 442, 176, 486, 626, 386, 291, 277, 638, 553, 568, 325]. multiway [458].

Natural [87, 123]. navigating [628]. navigation [168]. near [462]. Nearest [565, 211, 373]. Nearest-neighbor [565, 211, 373]. neighbor [373, 565, 211]. net [254]. net-effect [254]. Network [183, 666, 72, 625, 401, 639, 135, 507, 520, 221]. networks [283, 544, 324, 40, 615, 344, 243, 50, 187, 645, 493, 394, 468]. Neural [50, 599, 645]. neuroimaging [444, 442, 597]. Newton [138, 224]. Neyman [365]. Nomination [357, 297]. nonlinear [274, 281, 567]. Nonparametric [475, 496, 398, 309, 478, 383, 453, 476, 227, 516, 442, 618, 530, 477, 573, 588, 439]. nonstandard [514]. Nonstationarity [62, 282]. Norm [663]. normal [602]. normality [291]. normalization [577]. note [381]. Nuclear [663]. Number [659, 395, 461]. Numerical [174, 592].

Object [299, 526]. occurrences [456]. Ockham [76]. older [134]. OMARS [665]. omic [333]. omics [612, 550]. omnibus [403]. oncology [319]. One [665]. ontologies [165]. open [213, 597]. open-source [213]. Opioid [653]. opportunities [534]. optical [526]. Optimal [271, 110, 382, 74, 582]. Optimization [316, 2, 564, 586, 541, 135]. optimum [198]. Option [184]. order [244, 345, 555, 467]. Ordered [580]. orders [276]. Ordinal [655, 546, 352]. oriented [12, 299]. Orthogonal [98]. other [460, 589]. outcomes [625, 212]. outcomes-Understanding [625]. outlier [250]. outliers [250, 8]. output [481]. Overdose [653]. overlapping [456]. Overview [307, 537, 519, 288, 551, 444, 597, 355, 406].

P [326]. P-splines [326]. Package [651, 223]. packages [616]. paradigm [374]. paradox [77]. Parallel [505, 144]. parameter [538]. parameterizations [241]. parameters [291]. parametric [227, 530, 301]. Part [196, 540, 545, 418, 219, 234, 235, 543, 549]. Part-of-speech [196]. Partial [61, 245, 343, 367, 262, 325]. Particle [564]. Particularities [483]. Partition [656]. Partition-Based [656]. partitioning [37]. partly [612]. partnerships [632]. past [430]. path [375]. pathway [468]. pathways [256].

Pattern [94, 456, 243]. patterns [279]. Pearson [365]. pedagogy [338]. penalties [119, 300]. penalty [240, 333]. performance [295, 195, 614]. Periodic [369, 635]. Permutation [181, 418, 442, 276]. persistent [512]. personal [293]. perspective [457, 644, 230, 427, 507]. Perspectives [661, 332]. perturbation [501]. Pharmacokinetics [164]. Player [350]. plot [30]. Plots [250, 144, 179, 201, 254]. PLS [61]. point [234, 614]. Poisson [386, 563, 531]. Poisson-based [563]. polynomials [347]. population [233]. populations [250]. portfolio [516, 319]. portfolios [561]. positive [277]. positron [599, 213]. post [629]. post-quantum [629]. posterior [331]. potential [619]. power [414, 635, 589]. practical [351, 628]. practices [549]. Pratt [424]. precedence [405]. precedence-type [405]. precision [401]. Prediction [612, 563, 478, 528]. predictions [509]. presence [460, 393]. prices [184]. pricing [184]. Principal [96, 239, 198, 258, 281, 552]. prior [331, 574, 631, 434]. prior-free [331]. priors [460]. privacy [622, 605]. probabilistic [331]. probabilistically [583]. probabilities [350, 405, 456]. probability [341, 607, 601]. problems [417, 501, 442, 305, 582, 592]. procedure [502]. procedures [377, 28]. process [373]. processes [163, 412, 266, 91]. processing [483, 213, 87, 359]. Procrustes [102]. Programming [193, 337, 248]. Projection [26, 582, 61]. Projection-based [582]. projections [501]. promise [605]. propagation [527]. propensity [607]. properties [446, 639]. proportions [233]. protection [622]. proximity [78]. pseudo [395]. pseudo-random [395]. Pseudogrand [127]. pursuit [26].

Quadratic [337, 248, 421]. qualititative [366]. quality [132]. quantification [527]. quantiles [275]. quantitative [328]. quantum [629]. Quasi [342, 74, 637, 623]. Quasi-least [342]. quasi-Monte [637, 623]. quasi-optimal [74].

R [516, 221, 435, 339, 505, 616, 193, 543, 594, 27, 567, 259, 651, 237]. R-INLA [435]. Random [501, 659, 395, 200, 242]. Randomization [92, 311]. Randomized [55]. range [52]. Rank [108, 45, 227, 20]. rank-based [20]. Ranked [97]. ranking [377, 463]. rapid [349]. rates [647, 410]. Razor [76]. real [359]. realized [561]. recapture [255]. reciprocal [406]. recognition [267, 94, 526]. recommendation [256]. Record [106, 656, 313]. Record-Linkage [656]. records [534]. Recursive [37]. reduction [472, 329, 515, 390, 301, 450, 340]. refinements [446]. Regression [655, 559, 61, 240, 371, 583, 423, 476, 309, 519, 460, 336, 442, 563, 500, 518, 532, 330, 383, 11, 125, 553, 531, 406, 209, 362, 325, 546, 212, 540, 294]. regression-based [325]. regressions [345]. Regular [520]. Regularization [663, 417, 452, 153, 406, 178]. regulation [256]. Regulatory [661]. rejection [242]. Related [650]. Reliability [44, 580, 72, 160]. Rendering [64]. Reoptimization [204]. repair [410]. Reply [285]. Reporting [655]. representation [369, 379]. resampling [206]. research [288, 414, 347, 365, 627, 619]. resonance [327]. Response [665, 66, 166, 214, 16, 543, 549, 553, 424]. Responses [655]. restoration [14]. results [628]. Review [658, 661, 654, 471, 655, 665, 494, 650, 562, 656, 624, 435, 502, 423, 513, 564, 625, 476, 644, 580, 505, 556, 616, 482, 626, 612, 622, 390, 637, 623, 386, 585, 524, 541, 518, 601, 569, 587, 610, 260, 509, 615, 555, 421, 630, 577, 56, 308, 546, 627, 579]. Ridge [11, 125]. right [237]. Risk [653, 537, 451, 207, 410, 562]. risks [85, 525]. RNA [556, 628]. Roadmap [2, 165]. Robust [583, 329, 476, 343, 519, 643, 648, 355, 651, 250]. Robustness [140]. role [338, 295]. root [387]. rule [250, 39, 101]. rules [251].

safe [574]. safety [252]. Sample [172, 561, 403, 530]. sampler [547]. samples [250]. Sampling [516, 601, 271, 400, 16, 637, 623, 276, 361, 56, 97]. SAREV [556]. SAS [131]. SAS(R) [276]. satellite [168]. Scale [67, 324, 639]. scale-free [639]. scaled [552]. scaling [198, 400, 208]. scaling-down [400]. Scan [231]. Science [241, 632, 80, 21, 544, 374, 318, 451, 187, 597]. Science-based [241]. Sciences [17, 655]. Scientific [298]. score [607]. Scott [101]. screen [375]. Screening [665, 444, 147]. seamless [227]. searching [564]. second [555]. second-order [555]. Secure [659]. security [155, 107, 629]. Seismic [95]. seismology [83]. select [277]. selecting [237]. Selection [251, 443, 475, 286, 341, 326, 518, 574, 432, 377, 408, 434, 300, 319, 294, 596, 393]. selective [444, 355, 308]. Semi [272, 500]. semi- [500]. Semi-supervised [272]. Semiparametric [323, 591, 532]. sensor [243]. separable [422]. separation [460]. sequencing [556, 628]. Sequential [139, 586, 614, 428]. seriation [194]. series [416, 98, 253, 48, 483, 626, 433, 22, 512, 568, 498, 379]. server [410]. server-farm [410]. servers [410]. Service [12]. set [429, 97, 628, 84]. sets [258]. setting [229]. shape [479, 173]. shift [282]. shifted [68]. short [603]. Shrinkage [240, 496, 460, 121]. side [301]. side-effects [301]. Sieve [409]. sigma [234]. sigma-point [234]. signal [483]. signals [538]. significant [268]. SIMEST [319]. similarity [399]. Simpson [77]. SIMUGRAM [319]. Simulating [277]. Simulation [319, 43, 304, 264, 57, 49]. Simulation-based [319, 304]. simulations [338]. simulator [482]. simulator-models [482]. Simultaneous [385]. single [556]. single-cell [556]. singular [483, 222, 302]. singularity [635]. size [172]. sleep [269]. sliced [345]. small [455, 608, 436]. small-area [436]. Smooth [170]. smoother [220]. Smoothing [347, 591]. social [80, 634]. sociology [348]. Soft [465]. Software [160, 18, 255, 616, 151, 575, 232]. solution [273]. solve [274, 592]. Some [287, 265, 588]. something [237]. source [213]. South [19]. space [67, 508]. space-time [508]. spaces [514]. spanning [204]. Sparse [274, 565, 631, 461]. Spatial [435, 367, 191, 373, 565, 506, 349, 266, 539, 594, 575, 322]. Spatio [91, 436, 332]. Spatio-temporal [91, 436, 332]. spatiotemporal [506, 266, 241, 404]. Spectral [69]. spectrum [598, 483]. speech [196, 634].

Splines [119, 326, 75, 347]. Sports [655, 350, 606]. squares [61, 281, 73, 567, 342, 325, 245]. Stability [569]. stable [290, 557]. Stat [116]. Stata [130]. state [558]. state-of-the-art [558]. Stationary [498, 496]. STATIS [198]. Statistical [83, 51, 661, 581, 455, 442, 49, 182, 621, 195, 389, 430, 359, 237, 256, 18, 606, 636, 184, 287, 644, 203, 544, 599, 586, 556, 100, 561, 213, 305, 539, 585, 151, 509, 427, 507, 589, 614, 354, 17]. Statistics [21, 197, 137, 667, 173, 651, 162, 210, 3, 457, 25, 227, 565, 315, 208, 267, 318, 293, 451, 231, 346, 20, 355, 304, 124, 150, 332, 176, 215, 306, 1]. Statistics.com [236]. Steepest [128]. steganalysis [158]. steganography [158]. Step [665]. stereoscopic [179]. stereoscopy [270]. Stochastic [60, 307, 581, 367, 273]. stream [603]. Streaming [133]. structural [310]. structure [61, 266, 401, 520]. structured [89, 593, 534]. structures [422, 545]. studies [598]. study [513, 132, 612, 603, 269, 319, 589]. Sturges [39]. style [197]. Subgroup [604]. subject [559]. subpopulations [375]. subset [377]. such [589]. sufficient [515, 390, 450, 340]. suggestions [365]. summarization [641]. sums [500]. supervised [272]. Support [36, 153, 414, 420, 178]. Surface [665, 66]. Surrogate [661]. Survey [658, 591, 255, 324, 607, 629, 361, 60, 603, 344, 365, 570, 619, 527, 394, 592]. surveys [608]. survival [494, 289]. swarm [564]. symbolic [374]. symmetry [208]. symposia [318]. synthetic [622]. SYSTAT [79]. system [274, 367, 210]. Systematic [655, 616]. systems [168, 401, 23, 118].

tables [262, 461]. tagging [196]. Taguchi [177]. tail [416]. tailed [290, 247]. target [219, 234, 235]. targets [382]. tau [375]. tau-path [375]. taxonomy [546]. teaching [227]. Techniques [634, 591, 155, 203, 637, 623, 541, 538, 582]. technology [279]. temporal [91, 436, 332]. Tensor [472, 340]. Tessellation [303]. tessellations [109]. test [549]. testing [644, 422, 530, 593, 461, 492]. tests [403, 405, 561, 442, 387, 276, 530, 477, 588, 170]. text [351, 603, 526, 579]. their [606, 625, 227, 540]. theoretic [207]. theoretical [394]. Theory [662, 184, 292, 543, 549, 182, 617, 34, 142, 84]. There [265]. Thinking [374, 305]. Thompson [516]. Three [179, 198]. Three-dimensional [179]. Tikhonov [452]. Time [253, 433, 416, 48, 483, 626, 22, 508, 512, 525, 568, 359, 212, 498, 379]. time-to-event [525]. today [320]. Tolerance [602]. tomography [599, 213]. tool [237]. toolkit [232]. tools [606, 213, 249, 540]. top [375]. top- [375]. topic [351]. topology [183]. Total [73]. tour [261, 30, 127]. tours [558]. Tracing [125]. tracking [219, 234, 235]. Traditional [540]. traffic [377]. Transcriptomic [666]. Transductive [154]. Transfer [88]. transformation [227, 148]. translation [188]. transport [582]. Tree [658, 89, 429, 593, 204]. Tree-Based [658]. Tree-structured [89]. trees [280]. Trellis [112]. trends [634]. trial [139]. trials [172, 55]. trimming [643]. Truncation [662]. tutorial [182, 169, 176]. Tweedie [500]. two [636, 530]. two-sample [530]. type [82, 405]. unbiasedness [588]. uncertainties [628]. uncertainty [527]. Uncovering [279]. Understanding [625, 252]. Uniform [657]. uniqueness [448]. Unit [387]. univariate [471, 626, 263, 439]. unknown [409]. Unsupervised [618, 569]. Use [354, 279, 574]. used [416]. useful [301]. user [143]. Using [666, 304, 357, 350, 252, 159, 583, 423, 455, 608, 341, 326, 213, 618, 499, 276, 262, 434, 300, 347, 635, 645, 333]. Utilizing [653].

validation [432, 589]. value [222, 38, 302]. valued [48, 362]. values [203]. Vandervieren [250]. Variable [371, 286, 326, 336, 650, 300, 393, 448, 443, 434, 424]. variables [477, 461, 520]. variance [422, 561, 35, 462]. variances [308]. variates [15, 242, 277]. variation [436]. Vector [654, 36, 178, 153, 420, 493]. versus [625, 614]. Vertex [297, 357]. via [251, 599, 482, 433, 291]. view [401]. vision [197, 487]. ViSta [210]. VisTracks [81]. Visual [194, 210]. Visualization [232, 93, 81, 288, 499, 558]. Visualizing [335, 302]. Volatility [186, 562, 321, 314, 478]. Volume [503, 363, 9]. votes [354].

Washington [19]. wave [538]. Wavelet [120, 121]. wavelets [455]. way [198, 227]. Weibull [161]. Weighted [432, 334]. weighting [607, 361]. Where [315]. wide [589, 528]. Wilcoxon [588]. win [350]. WinBUGS [169, 434]. WINNER [65]. WIREs [19, 65, 1]. without [229]. wombling [349].

year [236]. Years [508].

Zero [540, 545]. Zero-inflated [540, 545]. zeros [389]. Zurbenko [86, 404].

References

Wegman:2009:IWC

[1] Edward J. Wegman, Yasmin H. Said, and David W. Scott. Introducing WIREs Computational Statistics. WIREs Computational Statistics, 1(1): 1–2, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Said:2009:RO

[2] Yasmin Said and Edward Wegman. Roadmap for optimization. WIREs Computational Statistics, 1(1):3–17, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bailey:2009:EMC

[3] David H. Bailey and Jonathan M. Borwein. Experimental mathematics and computational statistics. WIREs Computational Statistics, 1(1):12– 24, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108

(electronic). URL http://docserver.carma.newcastle.edu.au/1227/.

Kettenring:2009:MD

[4] Jon R. Kettenring. Massive datasets. WIREs Computational Statistics, 1(1):25–32, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Morgenthaler:2009:EDA

[5] Stephan Morgenthaler. Exploratory data analysis. WIREs Computational Statistics, 1(1):33–44, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2009:LM

[6] David W. Scott. The L2E method. WIREs Computational Statistics, 1 (1):45–51, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

DeJong:2009:EC

[7] Kenneth De Jong. Evolutionary computation. WIREs Computational Statistics, 1(1):52–56, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hadi:2009:DO

[8] Ali S. Hadi, A. H. M. Rahmatullah Imon, and Mark Werner. Detection of outliers. WIREs Computational Statistics, 1(1):57–70, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

VanAelst:2009:MVE

[9] Stefan Van Aelst and Peter Rousseeuw. Minimum volume ellipsoid. WIREs Computational Statistics, 1(1):71–82, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Duncan:2009:DMD

[10] George Duncan and Lynne Stokes. Data masking for disclosure limitation. WIREs Computational Statistics, 1(1):83–92, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McDonald:2009:RR

[11] Gary C. McDonald. Ridge regression. WIREs Computational Statistics, 1(1):93–100, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Laskey:2009:SOA

[12] Kathryn B. Laskey and Kenneth Laskey. Service oriented architecture. WIREs Computational Statistics, 1(1):101–105, July/August 2009. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Marchette:2009:FKD

[13] David Marchette. Filtered kernel density estimation. WIREs Computational Statistics, 1(1):106–109, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Qiu:2009:IR

[14] Peihua Qiu. Image restoration. WIREs Computational Statistics, 1(1): 110–113, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gentle:2009:AV

[15] James E. Gentle. Antithetic variates. WIREs Computational Statistics, 1(1):114–117, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hardwick:2009:ARA

[16] Janis Hardwick and Quentin F. Stout. Algorithms for response adaptive sampling designs. WIREs Computational Statistics, 1(1):118–122, July/ August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Berger:2009:SAM

[17] James Berger. Statistical and Applied Mathematical Sciences Institute. WIREs Computational Statistics, 1(1):123–127, July/August 2009. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

deLeeuw:2009:JSS

[18] Jan de Leeuw. Journal of statistical software. WIREs Computational Statistics, 1(1):128–129, July/August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wegman:2009:WBW

[19] Edward J. Wegman, Yasmin H. Said, and David W. Scott. WIREs between Washington, DC and Durban, South Africa. WIREs Computational Statistics, 1(2):131, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McKean:2009:CRB

[20] Joseph W. McKean, Jeff T. Terpstra, and John D. Kloke. Computational rank-based statistics. WIREs Computational Statistics, 1(2):132– 140, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Curran:2009:SFS

[21] James M. Curran. Statistics in forensic science. WIREs Computational Statistics, 1(2):141–156, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Politis:2009:FTS

[22] Dimitris N. Politis. Financial time series. WIREs Computational Statistics, 1(2):157–166, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wieczorek:2009:GIS

[23] William F. Wieczorek and Alan M. Delmerico. Geographic information systems. WIREs Computational Statistics, 1(2):167–186, September/ October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See erratum [118].

Brockwell:2009:A

[24] P. J. Brockwell. Autocovariance. WIREs Computational Statistics, 2(5): 187–198, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Conover:2009:DFM

[25] W. J. Conover. Distribution-free methods in statistics. WIREs Computational Statistics, 2(5):199–207, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jee:2009:PP

[26] J. Rodney Jee. Projection pursuit. WIREs Computational Statistics, 2 (5):208–215, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Murrell:2009:RG

[27] Paul Murrell. R graphics. WIREs Computational Statistics, 2(5):216– 220, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McKean:2009:DP

[28] Joseph W. McKean and Simon J. Sheather. Diagnostic procedures. WIREs Computational Statistics, 2(5):221–233, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gunst:2009:FFD

[29] Richard F. Gunst and Robert L. Mason. Fractional factorial design. WIREs Computational Statistics, 2(5):234–244, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Moustafa:2009:GTA

[30] Rida E. Moustafa and Ali S. Hadi. Grand tour and the Andrews plot. WIREs Computational Statistics, 2(5):245–250, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Higham:2009:CF

[31] Nicholas J. Higham. Cholesky factorization. WIREs Computational Statistics, 2(5):251–254, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Payne:2009:G

[32] Roger W. Payne. GenStat. WIREs Computational Statistics, 2(5):255– 258, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Abdi:2009:C

[33] Hervé Abdi. Centroids. WIREs Computational Statistics, 2(5):259–260, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Stine:2009:CT

[34] Robert A. Stine. Coding theory. WIREs Computational Statistics, 1 (3):261–270, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Muller:2009:AVC

[35] Keith E. Muller. Analysis of variance concepts and computations. WIREs Computational Statistics, 1(3):271–282, November/December 2009. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Mammone:2009:SVM

[36] Alessia Mammone, Marco Turchi, and Nello Cristianini. Support vector machines. WIREs Computational Statistics, 1(3):283–289, November/ December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hawkins:2009:RP

[37] Douglas M. Hawkins. Recursive partitioning. WIREs Computational Statistics, 1(3):290–295, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hubert:2009:BV

[38] Mia Hubert and Michiel Debruyne. Breakdown value. WIREs Computational Statistics, 1(3):296–302, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2009:SR

[39] David W. Scott. Sturges’ rule. WIREs Computational Statistics, 1(3):303– 306, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jensen:2009:BN

[40] Finn V. Jensen. Bayesian networks. WIREs Computational Statistics, 1 (3):307–315, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hopkins:2009:CAA

[41] Tim Hopkins. The Collected Algorithms of the ACM. WIREs Computational Statistics, 1(3):316–324, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chow:2009:BB

[42] Winston C. Chow. Brownian bridge. WIREs Computational Statistics, 1 (3):325–332, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2009:CAS

[43] Jim X. Chen. Continuous animation and simulation. WIREs Computational Statistics, 1(3):333–337, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wilson:2009:RC

[44] Simon P. Wilson. Reliability in the 21st century. WIREs Computational Statistics, 1(3):338–341, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wolfe:2009:RM

[45] Douglas A. Wolfe. Rank methods. WIREs Computational Statistics, 1 (3):342–347, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Nazareth:2009:CGM

[46] J. L. Nazareth. Conjugate gradient method. WIREs Computational Statistics, 1(3):348–353, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Piegorsch:2009:CI

[47] Walter W. Piegorsch and A. John Bailer. Combining information. WIREs Computational Statistics, 1(3):354–360, November/December 2009. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fokianos:2009:IVT

[48] Konstantinos Fokianos. Integer-valued time series. WIREs Computational Statistics, 1(3):361–364, November/December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Krutchkoff:2009:JSC

[49] Richard G. Krutchkoff. Journal of Statistical Computation and Simulation. WIREs Computational Statistics, 1(3):365–367, November/ December 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Titterington:2010:NN

[50] Michael Titterington. Neural networks. WIREs Computational Statistics, 2(1):1–8, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Banks:2010:SDM

[51] David L. Banks. Statistical data mining. WIREs Computational Statistics, 2(1):9–25, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Beran:2010:LRD

[52] Jan Beran. Long-range dependence. WIREs Computational Statistics, 2 (1):26–35, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hubert:2010:MCD

[53] Mia Hubert and Michiel Debruyne. Minimum covariance determinant. WIREs Computational Statistics, 2(1):36–43, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2010:H

[54] David W. Scott. Histogram. WIREs Computational Statistics, 2(1):44–48, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zheng:2010:RCT

[55] Lu Zheng and Marvin Zelen. Randomized clinical trials. WIREs Computational Statistics, 2(1):49–53, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tokdar:2010:ISR

[56] Surya T. Tokdar and Robert E. Kass. Importance sampling: a review. WIREs Computational Statistics, 2(1):54–60, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Thompson:2010:FSM

[57] James Thompson. Forward simulation models. WIREs Computational Statistics, 2(1):61–68, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Buhlmann:2010:B

[58] Peter Bühlmann and Bin Yu. Boosting. WIREs Computational Statistics, 2(1):69–74, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Soofi:2010:IM

[59] Ehsan S. Soofi, Huimin Zhao, and Derek L. Nazareth. Information measures. WIREs Computational Statistics, 2(1):75–86, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kushner:2010:SAS

[60] Harold Kushner. Stochastic approximation: a survey. WIREs Computational Statistics, 2(1):87–96, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Abdi:2010:PLS

[61] Hervé Abdi. Partial least squares regression and projection on latent structure regression (PLS regression). WIREs Computational Statistics, 2(1):97–106, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yang:2010:N

[62] Wei Yang and Igor Zurbenko. Nonstationarity. WIREs Computational Statistics, 2(1):107–115, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huo:2010:B

[63] Xiaoming Huo. Beamlets. WIREs Computational Statistics, 2(1):116–119, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2010:RCG

[64] Jim X. Chen. Rendering in computer graphics. WIREs Computational Statistics, 2(1):120–126, January/February 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wegman:2010:WW

[65] Edward J. Wegman, Yasmin H. Said, and David W. Scott. WIREs is a WINNER. WIREs Computational Statistics, 2(2):127, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Khuri:2010:RSM

[66] André I. Khuri and Siuli Mukhopadhyay. Response surface methodology. WIREs Computational Statistics, 2(2):128–149, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Holmstrom:2010:SSM

[67] Lasse Holmström. Scale space methods. WIREs Computational Statistics, 2(2):150–159, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2010:ASH

[68] David W. Scott. Averaged shifted histogram. WIREs Computational Statistics, 2(2):160–164, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fokianos:2010:SE

[69] Konstantinos Fokianos. Spectral estimation. WIREs Computational Statistics, 2(2):165–170, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Marchette:2010:CCC

[70] David Marchette. Class cover catch digraphs. WIREs Computational Statistics, 2(2):171–177, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gijbels:2010:CD

[71] Irène Gijbels. Censored data. WIREs Computational Statistics, 2(2):178– 188, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sahinoglu:2010:NRE

[72] Mehmet Sahinoglu and Benjamin Rice. Network reliability evaluation. WIREs Computational Statistics, 2(2):189–211, March/April 2010. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Markovsky:2010:TLS

[73] Ivan Markovsky, Diana M. Sima, and Sabine Van Huffel. Total least squares methods. WIREs Computational Statistics, 2(2):212–217, March/ April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cervone:2010:AQO

[74] Guido Cervone, Pasquale Franzese, and Allen P. K. Keesee. Algorithm quasi-optimal (AQ) learning. WIREs Computational Statistics, 2(2):218– 236, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gillies:2010:BS

[75] Duncan Gillies. B-splines. WIREs Computational Statistics, 2(2):237– 242, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lazar:2010:OR

[76] Nicole Lazar. Ockham’s Razor. WIREs Computational Statistics, 2(2): 243–246, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Alin:2010:SP

[77] Aylin Alin. Simpson’s paradox. WIREs Computational Statistics, 2(2): 247–250, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2010:BPM

[78] Angel R. Martinez. Bigram proximity matrix. WIREs Computational Statistics, 2(2):251–255, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wilkinson:2010:S

[79] Leland Wilkinson. SYSTAT. WIREs Computational Statistics, 2(2):256– 257, March/April 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cioffi-Revilla:2010:CSS

[80] Claudio Cioffi-Revilla. Computational social science. WIREs Computational Statistics, 2(3):259–271, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Eick:2010:GVV

[81] Stephen G. Eick. Geospatial visualization with VisTracks. WIREs Computational Statistics, 2(3):272–286, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Alnoshan:2010:DGT

[82] Abdullah A. Alnoshan and Shmuel Rotenstreich. Distributed grid-type architectures. WIREs Computational Statistics, 2(3):287–302, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anderson:2010:SMS

[83] Dale N. Anderson, Robert H. Shumway, Robert R. Blandford, and Steven R. Taylor. Statistical methods in seismology. WIREs Computational Statistics, 2(3):303–316, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zimmermann:2010:FST

[84] H.-J. Zimmermann. Fuzzy set theory. WIREs Computational Statistics, 2(3):317–332, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Klein:2010:CR

[85] John P. Klein. Competing risks. WIREs Computational Statistics, 2(3): 333–339, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yang:2010:KZF

[86] Wei Yang and Igor Zurbenko. Kolmogorov–Zurbenko filters. WIREs Computational Statistics, 2(3):340–351, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2010:NLP

[87] Angel R. Martinez. Natural language processing. WIREs Computational Statistics, 2(3):352–357, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Pollock:2010:TF

[88] Stephen Pollock. Transfer functions. WIREs Computational Statistics, 2(3):358–363, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Loh:2010:TSC

[89] Wei-Yin Loh. Tree-structured classifiers. WIREs Computational Statistics, 2(3):364–369, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Alin:2010:Ma

[90] Aylin Alin. Multicollinearity. WIREs Computational Statistics, 2(3):370– 374, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Harvill:2010:STP

[91] Jane L. Harvill. Spatio-temporal processes. WIREs Computational Statistics, 2(3):375–382, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Manly:2010:R

[92] Bryan F. J. Manly. Randomization. WIREs Computational Statistics, 2(3):383–386, May/June 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2010:IV

[93] Chaomei Chen. Information visualization. WIREs Computational Statistics, 2(4):387–403, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Holmstrom:2010:PR

[94] Lasse Holmström and Petri Koistinen. Pattern recognition. WIREs Computational Statistics, 2(4):404–413, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anderson:2010:SEI

[95] Dale N. Anderson, George E. Randall, Rodney W. Whitaker, Stephen J. Arrowsmith, Marie D. Arrowsmith, Deborah K. Fagan, Steven R. Taylor, Neil D. Selby, Frederick R. Schult, Gordon D. Kraft, and William R. Walter. Seismic event identification. WIREs Computational Statistics, 2 (4):414–432, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Abdi:2010:PCA

[96] Hervé Abdi and Lynne J. Williams. Principal component analysis. WIREs Computational Statistics, 2(4):433–459, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wolfe:2010:RSS

[97] Douglas A. Wolfe. Ranked set sampling. WIREs Computational Statistics, 2(4):460–466, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Efromovich:2010:OSD

[98] Sam Efromovich. Orthogonal series density estimation. WIREs Computational Statistics, 2(4):467–476, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lee:2010:CD

[99] Alan Lee. Circular data. WIREs Computational Statistics, 2(4):477–486, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fraser:2010:BIA

[100] D. A. S. Fraser. Bayesian inference: an approach to statistical inference. WIREs Computational Statistics, 2(4):487–496, July/August 2010. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2010:SR

[101] David W. Scott. Scott’s rule. WIREs Computational Statistics, 2(4):497– 502, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gower:2010:PM

[102] John C. Gower. Procrustes methods. WIREs Computational Statistics, 2 (4):503–508, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Verboven:2010:MLL

[103] Sabine Verboven and Mia Hubert. MATLAB library LIBRA. WIREs Computational Statistics, 2(4):509–515, July/August 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Reid:2010:LI

[104] Nancy Reid. Likelihood inference. WIREs Computational Statistics, 2 (5):517–525, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Holmstrom:2010:B

[105] Lasse Holmström. BSiZer. WIREs Computational Statistics, 2(5):526– 534, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Herzog:2010:RL

[106] Thomas H. Herzog, Fritz Scheuren, and William E. Winkler. Record linkage. WIREs Computational Statistics, 2(5):535–543, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gollmann:2010:CS

[107] Dieter Gollmann. Computer security. WIREs Computational Statistics, 2(5):544–554, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lin:2010:RAM

[108] Shili Lin. Rank aggregation methods. WIREs Computational Statistics, 2(5):555–570, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shores:2010:BDT

[109] Roger W. Shores and Edward J. Wegman. Bounds on Delaunay tessellations. WIREs Computational Statistics, 2(5):571–580, September/ October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fedorov:2010:OED

[110] Valerii Fedorov. Optimal experimental design. WIREs Computational Statistics, 2(5):581–589, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gijbels:2010:L

[111] Irène Gijbels and Ilaria Prosdocimi. Loess. WIREs Computational Statistics, 2(5):590–599, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Robbins:2010:TD

[112] Naomi B. Robbins. Trellis display. WIREs Computational Statistics, 2 (5):600–605, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Theus:2010:M

[113] Martin Theus. Mondrian. WIREs Computational Statistics, 2(5):606– 612, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Greenacre:2010:CA

[114] Michael J. Greenacre. Correspondence analysis. WIREs Computational Statistics, 2(5):613–619, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Johnson:2010:CE

[115] Dallas E. Johnson. Crossover experiments. WIREs Computational Statistics, 2(5):620–625, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tierney:2010:LS

[116] Luke Tierney. Lisp-Stat. WIREs Computational Statistics, 2(5):626–630, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shyamalkumar:2010:AM

[117] Nariankadu Shyamalkumar. Actuarial method. WIREs Computational Statistics, 2(5):631–634, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wieczorek:2010:EGI

[118] William F. Wieczorek and Alan M. Delmerico. Erratum: Geographic information systems. WIREs Computational Statistics, 2(5):635–636, September/October 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See [23].

Eilers:2010:SKP

[119] Paul H. C. Eilers and Brian D. Marx. Splines, knots, and penalties. WIREs Computational Statistics, 2(6):637–653, November/December 2010. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fryzlewicz:2010:WM

[120] Piotr Fryzlewicz. Wavelet methods. WIREs Computational Statistics, 2 (6):654–667, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huerta:2010:BWS

[121] Gabriel Huerta. Bayesian wavelet shrinkage. WIREs Computational Statistics, 2(6):668–672, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wilkinson:2010:GG

[122] Leland Wilkinson. The grammar of graphics. WIREs Computational Statistics, 2(6):673–677, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wegman:2010:NHC

[123] Edward J. Wegman and Yasmin H. Said. Natural homogeneous coordinates. WIREs Computational Statistics, 2(6):678–685, November/ December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Vos:2010:GS

[124] Paul W. Vos and Paul Marriott. Geometry in statistics. WIREs Computational Statistics, 2(6):686–694, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McDonald:2010:TRR

[125] Gary C. McDonald. Tracing ridge regression coefficients. WIREs Computational Statistics, 2(6):695–703, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Theus:2010:B

[126] Martin Theus. Brushing. WIREs Computational Statistics, 2(6):704–710, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Moustafa:2010:PT

[127] Rida E. Moustafa. Pseudogrand tour. WIREs Computational Statistics, 2(6):711–718, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Meza:2010:SD

[128] Juan C. Meza. Steepest descent. WIREs Computational Statistics, 2 (6):719–722, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Alin:2010:Mb

[129] Aylin Alin. Minitab. WIREs Computational Statistics, 2(6):723–727, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gutierrez:2010:S

[130] Roberto G. Gutierrez. Stata. WIREs Computational Statistics, 2(6):728– 733, November/December 2010. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rodriguez:2011:S

[131] Robert N. Rodriguez. SAS. WIREs Computational Statistics, 3(1):1–11, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Herzog:2011:DQC

[132] Thomas N. Herzog, Fritz J. Scheuren, and William E. Winkler. Data quality: a case study. WIREs Computational Statistics, 3(1):12–21, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Szewczyk:2011:SD

[133] William Szewczyk. Streaming data. WIREs Computational Statistics, 3 (1):22–29, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Caudle:2011:DOD

[134] Kyle A. Caudle, Gary O. Fowler, Leah R. Jager, and David M. Ruth. Discounting older data. WIREs Computational Statistics, 3(1):30–33, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rezazad:2011:CNO

[135] Hadi Rezazad. Computer network optimization. WIREs Computational Statistics, 3(1):34–46, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sahinoglu:2011:CC

[136] Mehmet Sahinoglu and Luis Cueva-Parra. CLOUD computing. WIREs Computational Statistics, 3(1):47–68, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2011:CSM

[137] Wendy L. Martinez. Computational statistics in MATLAB(R). WIREs Computational Statistics, 3(1):69–74, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Meza:2011:NM

[138] Juan C. Meza. Newton’s method. WIREs Computational Statistics, 3(1): 75–78, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dong:2011:SDT

[139] Ting Dong and Liansheng Tang. Sequential diagnostic trial designs. WIREs Computational Statistics, 3(1):79–83, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Morgenthaler:2011:R

[140] Stephan Morgenthaler. Robustness. WIREs Computational Statistics, 3 (2):85–94, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kettenring:2011:CHD

[141] Jon R. Kettenring. Coping with high dimensionality in massive datasets. WIREs Computational Statistics, 3(2):95–103, March/April 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wegman:2011:CTD

[142] Edward Wegman and Yasmin Said. Color theory and design. WIREs Computational Statistics, 3(2):104–117, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2011:GUI

[143] Wendy L. Martinez. Graphical user interfaces. WIREs Computational Statistics, 3(2):119–133, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Moustafa:2011:PCP

[144] Rida E. Moustafa. Parallel coordinate and parallel coordinate density plots. WIREs Computational Statistics, 3(2):134–148, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chow:2011:FFB

[145] Winston C. Chow. Fractal (fractional) Brownian motion. WIREs Computational Statistics, 3(2):149–162, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Krantz:2011:HA

[146] Steven G. Krantz. Harmonic analysis. WIREs Computational Statistics, 3(2):163–167, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tarter:2011:MFD

[147] Michael E. Tarter. Model-free data screening and cleaning. WIREs Computational Statistics, 3(2):168–176, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2011:BMT

[148] David W. Scott. Box–Muller transformation. WIREs Computational Statistics, 3(2):177–179, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wickham:2011:G

[149] Hadley Wickham. ggplot2. WIREs Computational Statistics, 3(2):180– 185, March/April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wegman:2011:MLD

[150] Edward J. Wegman, Yasmin H. Said, and David W. Scott. Machine learning, data mining, and computational statistics applications. WIREs Computational Statistics, 3(3):187, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jones:2011:JSD

[151] Bradley Jones and John Sall. JMP statistical discovery software. WIREs Computational Statistics, 3(3):188–194, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Alpaydin:2011:ML

[152] Ethem Alpaydin. Machine learning. WIREs Computational Statistics, 3(3):195–203, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Reeves:2011:SVM

[153] D. M. Reeves and G. M. Jacyna. Support vector machine regularization. WIREs Computational Statistics, 3(3):204–215, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See erratum [178].

Rogers:2011:TCM

[154] James Rogers. Transductive confidence machines. WIREs Computational Statistics, 3(5):216–220, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Al-Shawi:2011:DMT

[155] Amany Al-Shawi. Data mining techniques for information security applications. WIREs Computational Statistics, 3(5):221–229, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Higham:2011:GE

[156] Nicholas J. Higham. Gaussian elimination. WIREs Computational Statistics, 3(5):230–238, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Stallings:2011:C

[157] William Stallings. Ciphers. WIREs Computational Statistics, 3(5):239– 250, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bachrach:2011:ISS

[158] Mayra Bachrach and Frank Y. Shih. Image steganography and steganalysis. WIREs Computational Statistics, 3(5):251–259, May/June 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Barrett:2011:CUA

[159] Timothy J. Barrett. Computations using analysis of covariance. WIREs Computational Statistics, 3(5):260–268, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Soyer:2011:SR

[160] Refik Soyer. Software reliability. WIREs Computational Statistics, 3(5): 269–281, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lai:2011:WD

[161] C. D. Lai, D. N. P. Murthy, and M. Xie. Weibull distributions. WIREs Computational Statistics, 3(5):282–287, May/June 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tsay:2011:SF

[162] Ruey S. Tsay. Statistics in finance. WIREs Computational Statistics, 3 (4):289–315, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Brockwell:2011:AP

[163] P. J. Brockwell. Autoregressive processes. WIREs Computational Statistics, 3(4):316–331, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bonate:2011:P

[164] Peter L. Bonate. Pharmacokinetics. WIREs Computational Statistics, 3 (4):332–342, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

An:2011:ABM

[165] Gary An and Scott Christley. Agent-based modeling and biomedical ontologies: a roadmap. WIREs Computational Statistics, 3(4):343–356, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shore:2011:RMM

[166] Haim Shore. Response modeling methodology. WIREs Computational Statistics, 3(4):357–372, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Moustafa:2011:AC

[167] Rida E. Moustafa. Andrews curves. WIREs Computational Statistics, 3 (4):373–382, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Grewal:2011:GNS

[168] Mohinder S. Grewal. Global navigation satellite systems. WIREs Computational Statistics, 3(4):383–384, July/August 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lykou:2011:WT

[169] Anastasia Lykou and Ioannis Ntzoufras. WinBUGS: a tutorial. WIREs Computational Statistics, 3(5):385–396, September/October 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rayner:2011:STG

[170] J. C. W. Rayner, O. Thas, and D. J. Best. Smooth tests of goodness of fit. WIREs Computational Statistics, 3(5):397–406, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Neuhaus:2011:GLM

[171] John Neuhaus and Charles McCulloch. Generalized linear models. WIREs Computational Statistics, 3(5):407–413, September/October 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chow:2011:SSC

[172] Shein-Chung Chow. Sample size calculations for clinical trials. WIREs Computational Statistics, 3(5):414–427, September/October 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Small:2011:SS

[173] Christopher G. Small. Statistics of shape. WIREs Computational Statistics, 3(5):428–433, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Brandimarte:2011:NA

[174] Paolo Brandimarte. Numerical analysis. WIREs Computational Statistics, 3(5):434–449, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Reiter:2011:DC

[175] Jerome P. Reiter. Data confidentiality. WIREs Computational Statistics, 3(5):450–456, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yu:2011:MST

[176] Q. Yu, W. Tang, J. Kowalski, and X. M. Tu. Multivariate U-statistics: a tutorial with applications. WIREs Computational Statistics, 3(5):457– 471, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Mitra:2011:TM

[177] Amitava Mitra. The Taguchi method. WIREs Computational Statistics, 3(5):472–480, September/October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Reeves:2011:ESV

[178] D. M. Reeves and G. M. Jacyna. Erratum: Support vector machine regularization. WIREs Computational Statistics, 3(5):481, September/ October 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See [153].

Symanzik:2011:TDS

[179] Jürgen Symanzik. Three-dimensional stereoscopic plots. WIREs Computational Statistics, 3(6):483–496, November/December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hesterberg:2011:B

[180] Tim Hesterberg. Bootstrap. WIREs Computational Statistics, 3(6):497– 526, November/December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Berry:2011:PM

[181] Kenneth J. Berry, Janis E. Johnston, and Paul W. Mielke Jr. Permutation methods. WIREs Computational Statistics, 3(6):527–542, November/ December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kulkarni:2011:SLT

[182] Sanjeev R. Kulkarni and Gilbert Harman. Statistical learning theory: a tutorial. WIREs Computational Statistics, 3(6):543–556, November/ December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kincaid:2011:NTM

[183] Rex K. Kincaid and David J. Phillips. Network topology measures. WIREs Computational Statistics, 3(6):557–565, November/December 2011. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2011:OPP

[184] Ling Chen, Tze Leung Lai, and Tiong Wee Lim. Option prices and pricing theory: combining financial mathematics with statistical modeling. WIREs Computational Statistics, 3(6):566–576, November/December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dey:2011:FD

[185] Aloke Dey. Factorial designs. WIREs Computational Statistics, 3(6):577– 581, November/December 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Laurent:2012:VFE

[186] Sébastien Laurent and Francesco Violante. Volatility forecasts evaluation and comparison. WIREs Computational Statistics, 4(1):1–12, January/ February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Vivar:2012:MNC

[187] Juan C. Vivar and David Banks. Models for networks: a cross-disciplinary science. WIREs Computational Statistics, 4(1):13–27, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Marchette:2012:IT

[188] David Marchette. Implicit translation. WIREs Computational Statistics, 4 (1):28–34, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chi:2012:MM

[189] Yueh-Yun Chi. Multivariate methods. WIREs Computational Statistics, 4 (1):35–47, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kroese:2012:MCM

[190] Dirk P. Kroese and Reuven Y. Rubinstein. Monte Carlo methods. WIREs Computational Statistics, 4(1):48–58, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Banerjee:2012:BML

[191] Sudipto Banerjee and Montserrat Fuentes. Bayesian modeling for large spatial datasets. WIREs Computational Statistics, 4(1):59–66, January/ February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wieczorek:2012:CIA

[192] William F. Wieczorek, Alan M. Delmerico, Peter A. Rogerson, and David W. S. Wong. Clusters in irregular areas and lattices. WIREs Computational Statistics, 4(1):67–74, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Horgan:2012:PR

[193] Jane M. Horgan. Programming in R. WIREs Computational Statistics, 4 (1):75–84, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liiv:2012:VME

[194] Innar Liiv, Rain Opik, Jaan Ubi, and John Stasko. Visual matrix explorer for collaborative seriation. WIREs Computational Statistics, 4(1):85–97, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Patil:2012:SMC

[195] Shruti Patil and David J. Lilja. Statistical methods for computer performance evaluation. WIREs Computational Statistics, 4(1):98–106, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2012:PST

[196] Angel R. Martinez. Part-of-speech tagging. WIREs Computational Statistics, 4(1):107–113, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Graham:2012:SVA

[197] Daniel J. Graham, James M. Hughes, Helmut Leder, and Daniel N. Rockmore. Statistics, vision, and the analysis of artistic style. WIREs Computational Statistics, 4(2):115–123, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Abdi:2012:SDO

[198] Hervé Abdi, Lynne J. Williams, Domininique Valentin, and Mohammed Bennani-Dosse. STATIS and DISTATIS: optimum multitable principal component analysis and three way metric multidimensional scaling. WIREs Computational Statistics, 4(2):124–167, March/April 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tecuci:2012:AI

[199] Gheorghe Tecuci. Artificial intelligence. WIREs Computational Statistics, 4(2):168–180, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gurka:2012:FRE

[200] Matthew J. Gurka, George A. Kelley, and Lloyd J. Edwards. Fixed and random effects models. WIREs Computational Statistics, 4(2):181–190, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Theus:2012:MP

[201] Martin Theus. Mosaic plots. WIREs Computational Statistics, 4(2):191– 198, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Neath:2012:BIC

[202] Andrew A. Neath and Joseph E. Cavanaugh. The Bayesian information criterion: background, derivation, and applications. WIREs Computational Statistics, 4(2):199–203, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

DeWaal:2012:ESD

[203] Ton De Waal, Jeroen Pannekoek, and Sander Scholtus. The editing of statistical data: methods and techniques for the efficient detection and correction of errors and missing values. WIREs Computational Statistics, 4(2):204–210, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Paschos:2012:RMS

[204] Stratos A. Paschos and Vangelis Th. Paschos. Reoptimization of the minimum spanning tree. WIREs Computational Statistics, 4(2):211–217, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

vonEye:2012:LLM

[205] Alexander von Eye, Eun-Young Mun, and Patrick Mair. Log-linear modeling. WIREs Computational Statistics, 4(2):218–223, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chernick:2012:JRM

[206] Michael R. Chernick. The jackknife: a resampling method with connections to the bootstrap. WIREs Computational Statistics, 4(2):224–226, March/April 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sahinoglu:2012:GTC

[207] Mehmet Sahinoglu, Luis Cueva-Parra, and David Ang. Game-theoretic computing in risk analysis. WIREs Computational Statistics, 4(3):227– 248, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Frame:2012:FMG

[208] Michael Frame and William Martino. Fractal and multifractal geometry: scaling symmetry and statistics. WIREs Computational Statistics, 4(3): 249–274, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Su:2012:LR

[209] Xiaogang Su, Xin Yan, and Chih-Ling Tsai. Linear regression. WIREs Computational Statistics, 4(3):275–294, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Valero-Mora:2012:HVV

[210] Pedro Valero-Mora, Rubén D. Ledesma, and Michael Friendly. The history of ViSta: The Visual Statistics System. WIREs Computational Statistics, 4(3):295–306, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sutton:2012:NNM

[211] Clifton Sutton. Nearest-neighbor methods. WIREs Computational Statistics, 4(3):307–309, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Walters:2012:ATE

[212] Stephen J. Walters. Analyzing time to event outcomes with a Cox regression model. WIREs Computational Statistics, 4(3):310–315, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hawe:2012:KAD

[213] David Hawe, Francisco R. Hernández Fernández, Liam O’Suilleabháin, Jian Huang, Eric Wolsztynski, and Finbarr O’Sullivan. Kinetic analysis of dynamic positron emission tomography data using open-source image processing and statistical inference tools. WIREs Computational Statistics, 4(3):316–322, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shore:2012:ERM

[214] Haim Shore. Estimating response modeling methodology models. WIREs Computational Statistics, 4(3):323–333, May/June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hardle:2012:CSJ

[215] Wolfgang Karl Härdle, Yuichi Mori, and Jürgen Symanzik. Computational Statistics (journal). WIREs Computational Statistics, 4(3):334–339, May/ June 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2012:IIa

[216] Anonymous. Issue information. WIREs Computational Statistics, 4(4): I–II, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Stahl:2012:MBC

[217] Daniel Stahl and Hannah Sallis. Model-based cluster analysis. WIREs Computational Statistics, 4(4):341–358, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

El-Askary:2012:CMC

[218] Hesham El-Askary, Mohamed Allali, Cyril Rakovski, Anup Prasad, Menas Kafatos, and Daniele Struppa. Computational methods for climate data. WIREs Computational Statistics, 4(4):359–374, July/August 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Haug:2012:BETa

[219] A. J. Haug. Bayesian estimation for target tracking: Part I, general concepts. WIREs Computational Statistics, 4(4):375–383, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Klemela:2012:S

[220] Jussi Klemelä. Bin smoother. WIREs Computational Statistics, 4(4):384– 393, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hornik:2012:CRA

[221] Kurt Hornik. The Comprehensive R Archive Network. WIREs Computational Statistics, 4(4):394–398, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Greenacre:2012:BJS

[222] Michael J. Greenacre. Biplots: the joy of singular value decomposition. WIREs Computational Statistics, 4(4):399–406, July/August 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Velleman:2012:DIP

[223] Paul F. Velleman. DataDesk: an interactive package for data exploration, display, model building, and data analysis. WIREs Computational Statistics, 4(4):407–414, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2012:GNM

[224] Yong Wang. Gauss–Newton method. WIREs Computational Statistics, 4 (4):415–420, July/August 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2012:IIb

[225] Anonymous. Issue information. WIREs Computational Statistics, 4(5): I–II, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McLachlan:2012:DA

[226] Geoffrey J. McLachlan. Discriminant analysis. WIREs Computational Statistics, 4(5):421–431, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Conover:2012:RTE

[227] W. Jay Conover. The rank transformation — an easy and intuitive way to connect many nonparametric methods to their parametric counterparts for seamless teaching introductory statistics courses. WIREs Computational Statistics, 4(5):432–438, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Izenman:2012:IML

[228] Alan Julian Izenman. Introduction to manifold learning. WIREs Computational Statistics, 4(5):439–446, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Goldman:2012:MGS

[229] William M. Goldman. Manifolds: a global setting for geometry and analysis, with or without coordinates. WIREs Computational Statistics, 4 (5):447–454, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lodhi:2012:CBP

[230] Huma Lodhi. Computational biology perspective: kernel methods and deep learning. WIREs Computational Statistics, 4(5):455–465, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Marchette:2012:SSG

[231] David Marchette. Scan statistics on graphs. WIREs Computational Statistics, 4(5):466–473, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wills:2012:VTS

[232] Graham Wills. Visualization toolkit software. WIREs Computational Statistics, 4(5):474–481, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Okafor:2012:BMI

[233] R. O. Okafor and U. A. Mbata. A Bayesian model for inference on population proportions. WIREs Computational Statistics, 4(5):482–488, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Haug:2012:BETb

[234] A. J. Haug. Bayesian estimation for target tracking: Part II, the Gaussian sigma-point Kalman filters. WIREs Computational Statistics, 4(5):489– 497, September/October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Haug:2012:BETc

[235] A. J. Haug. Bayesian estimation for target tracking, Part III: Monte Carlo filters. WIREs Computational Statistics, 4(5):498–512, September/ October 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bruce:2012:SCY

[236] Peter Bruce. Statistics.com at the 10-year mark. WIREs Computational Statistics, 4(6):513–517, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wilson:2012:SCR

[237] Jason Wilson. Statistical computing with R: selecting the right tool for the job — R Commander or something else? WIREs Computational Statistics, 4(6):518–526, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liu:2012:FME

[238] Ziyue Liu and Wensheng Guo. Functional mixed effects models. WIREs Computational Statistics, 4(6):527–534, November/December 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Billard:2012:PCA

[239] L. Billard and J. Le-Rademacher. Principal component analysis for interval data. WIREs Computational Statistics, 4(6):535–540, November/ December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ahmed:2012:SAP

[240] S. Ejaz Ahmed and S. M. Enayetur Raheem. Shrinkage and absolute penalty estimation in linear regression models. WIREs Computational Statistics, 4(6):541–553, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Leeds:2012:SBP

[241] William B. Leeds and Christopher K. Wikle. Science-based parameterizations for dynamical spatiotemporal models. WIREs Computational Statistics, 4(6):554–560, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hall:2012:DRA

[242] Timothy Hall. Dominated rejection algorithms for generating random variates. WIREs Computational Statistics, 4(6):561–564, November/ December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shine:2012:AAP

[243] James A. Shine and James E. Gentle. Alarm activation, pattern discovery, and anomaly detection in sensor networks. WIREs Computational Statistics, 4(6):565–570, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chan:2012:HOA

[244] N. H. Chan and S. C. P. Yam. Higher-order asymptotics in finance. WIREs Computational Statistics, 4(6):571–587, November/December 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Vinzi:2013:PLS

[245] Vincenzo Esposito Vinzi and Giorgio Russolillo. Partial least squares algorithms and methods. WIREs Computational Statistics, 5(1):1–19, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rassler:2013:I

[246] Susanne Rässler, Donald B. Rubin, and Elizabeth R. Zell. Imputation. WIREs Computational Statistics, 5(1):20–29, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rojo:2013:HTD

[247] Javier Rojo. Heavy-tailed densities. WIREs Computational Statistics, 5 (1):30–40, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Xu:2013:QPA

[248] Jie Xu and J. Brian Gray. Quadratic programming algorithms for ensemble models. WIREs Computational Statistics, 5(1):41–47, January/ February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Loy:2013:DTH

[249] Adam Loy and Heike Hofmann. Diagnostic tools for hierarchical linear models. WIREs Computational Statistics, 5(1):48–61, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bradstreet:2013:HDP

[250] Thomas E. Bradstreet, Amy Gillespie, Shuping Zhang, and Xingshu Zhu. Hanging dot plots and identifying outliers in samples from asymmetrical populations with Hubert and Vandervieren’s robust outlier rule. WIREs Computational Statistics, 5(1):62–67, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Changpetch:2013:SML

[251] Pannapa Changpetch and Dennis K. J. Lin. Selection of multinomial logit models via association rules analysis. WIREs Computational Statistics, 5 (1):68–77, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anziano:2013:EUS

[252] Richard Anziano and Robert Gordon. Examples of understanding safety data using graphical methods. WIREs Computational Statistics, 5(1): 78–95, January/February 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ensor:2013:TSF

[253] Katherine Bennett Ensor. Time series factor models. WIREs Computational Statistics, 5(2):97–104, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2013:NDN

[254] Xin Zhang. New developments for net-effect plots. WIREs Computational Statistics, 5(2):105–113, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bunge:2013:SSF

[255] John A. Bunge. A survey of software for fitting capture-recapture models. WIREs Computational Statistics, 5(2):114–120, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yoo:2013:SME

[256] Changwon Yoo. Statistical methods in experimentation recommendation models for discovering gene regulation pathways. WIREs Computational Statistics, 5(2):121–134, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Melnykov:2013:CMB

[257] Volodymyr Melnykov. Challenges in model-based clustering. WIREs Computational Statistics, 5(2):135–148, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Abdi:2013:MFA

[258] Hervé Abdi, Lynne J. Williams, and Domininique Valentin. Multiple factor analysis: principal component analysis for multitable and multiblock data sets. WIREs Computational Statistics, 5(2):149–179, March/April 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Runnalls:2013:CER

[259] Andrew Runnalls. CXXR: an extensible R interpreter. WIREs Computational Statistics, 5(3):181–189, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Mai:2013:RDA

[260] Qing Mai. A review of discriminant analysis in high dimensions. WIREs Computational Statistics, 5(3):190–197, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2013:IGT

[261] Wendy L. Martinez. Image grand tour. WIREs Computational Statistics, 5(3):198–206, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Karwa:2013:CIG

[262] Vishesh Karwa and Aleksandra Slavkovic. Conditional inference given partial information in contingency tables using Markov bases. WIREs Computational Statistics, 5(3):207–218, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lee:2013:MGF

[263] Carl Lee, Felix Famoye, and Ayman Y. Alzaatreh. Methods for generating families of univariate continuous distributions in the recent decades. WIREs Computational Statistics, 5(3):219–238, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sahinoglu:2013:MSE

[264] Mehmet Sahinoglu. Modeling and simulation in engineering. WIREs Computational Statistics, 5(3):239–266, May/June 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Box:2013:TGB

[265] George E. P. Box. “There goes Bill!”: Bill Hunter and some ideas on experimental design. WIREs Computational Statistics, 5(4):267–278, July/ August 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Guttorp:2013:CSS

[266] Peter Guttorp and Alexandra M. Schmidt. Covariance structure of spatial and spatiotemporal processes. WIREs Computational Statistics, 5(4):279– 287, July/August 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Givens:2013:BFR

[267] Geof H. Givens, J. Ross Beveridge, Yui Man Lui, David S. Bolme, Bruce A. Draper, and P. Jonathon Phillips. Biometric face recognition: from classical statistics to future challenges. WIREs Computational Statistics, 5 (4):288–308, July/August 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bair:2013:ISF

[268] Eric Bair. Identification of significant features in DNA microarray data. WIREs Computational Statistics, 5(4):309–325, July/August 2013. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Piryatinska:2013:ESS

[269] Alexandra Piryatinska. Electroencephalogram-sleep study. WIREs Computational Statistics, 5(4):326–333, July/August 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

King:2013:BHS

[270] R. Duane King. A brief history of stereoscopy. WIREs Computational Statistics, 5(4):334–340, July/August 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2013:OMC

[271] Ting-Li Chen. Optimal Markov chain Monte Carlo sampling. WIREs Computational Statistics, 5(5):341–348, September/October 2013. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bair:2013:SSC

[272] Eric Bair. Semi-supervised clustering methods. WIREs Computational Statistics, 5(5):349–361, September/October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sauer:2013:CSS

[273] Timothy Sauer. Computational solution of stochastic differential equations. WIREs Computational Statistics, 5(5):362–371, September/ October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hossain:2013:SMC

[274] Shahadat Hossain and Trond Steihaug. Sparse matrix computations with application to solve system of nonlinear equations. WIREs Computational Statistics, 5(5):372–386, September/October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ren:2013:MEB

[275] Haobo Ren, Weining Shen, Richard Wu, and Yuhwen Soo. Moment estimation based on quantiles. WIREs Computational Statistics, 5(5):387– 390, September/October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

J:2013:BPT

[276] John D. (“J. D.”) Opdyke. Bootstraps, permutation tests, and sampling orders of magnitude faster using SAS(R). WIREs Computational Statistics, 5(5):391–405, September/October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Minhajuddin:2013:SSF

[277] Abu T. Minhajuddin and Sergio F. Juarez. Simulating select families of multivariate discrete variates with positive correlations. WIREs Computational Statistics, 5(5):406–414, September/October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ng:2013:RDE

[278] Shu-Kay Ng. Recent developments in expectation-maximization methods for analyzing complex data. WIREs Computational Statistics, 5(6):415– 431, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hung:2013:UPT

[279] Man Hung, Jillian Conrad, Shirley D. Hon, Christine Cheng, Jeremy D. Franklin, and Philip Tang. Uncovering patterns of technology use in consumer health informatics. WIREs Computational Statistics, 5(6):432–447, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

deVille:2013:DT

[280] Barry de Ville. Decision trees. WIREs Computational Statistics, 5(6):448– 455, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kuroda:2013:ALS

[281] Masahiro Kuroda, Yuichi Mori, Iizuka Masaya, and Michio Sakakihara. Alternating least squares in nonlinear principal components. WIREs Computational Statistics, 5(6):456–464, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sugiyama:2013:LUN

[282] Masashi Sugiyama, Makoto Yamada, and Marthinus Christoffel du Plessis. Learning under nonstationarity: covariate shift and class-balance change. WIREs Computational Statistics, 5(6):465–477, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2013:GEI

[283] Shyh-Huei Chen, Edward H. Ip, and Yuchung J. Wang. Gibbs ensembles for incompatible dependency networks. WIREs Computational Statistics, 5(6):478–485, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Brunner:2013:LE

[284] Edgar Brunner and Madan L. Puri. Letter to the Editor. WIREs Computational Statistics, 5(6):486–487, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Conover:2013:R

[285] W. J. Conover. Reply. WIREs Computational Statistics, 5(6):487–488, November/December 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2014:VSL

[286] Yuqi Chen, Pang Du, and Yuedong Wang. Variable selection in linear models. WIREs Computational Statistics, 6(1):1–9, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2014:SRS

[287] Shuo Chen, Edward Grant, Tong Tong Wu, and F. DuBois Bowman. Some recent statistical learning methods for longitudinal high-dimensional data. WIREs Computational Statistics, 6(1):10–18, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gan:2014:DVO

[288] Qihong Gan, Min Zhu, Mingzhao Li, Ting Liang, Yu Cao, and Baoyao Zhou. Document visualization: an overview of current research. WIREs Computational Statistics, 6(1):19–36, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liu:2014:HCA

[289] Yajun Liu, Dongchu Sun, and Chong Z. He. A hierarchical conditional autoregressive model for colorectal cancer survival data. WIREs Computational Statistics, 6(1):37–44, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Nolan:2014:FMH

[290] John P. Nolan. Financial modeling with heavy-tailed stable distributions. WIREs Computational Statistics, 6(1):45–55, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jamshidian:2014:EMD

[291] Mortaza Jamshidian and Ke-Hai Yuan. Examining missing data mechanisms via homogeneity of parameters, homogeneity of distributions, and multivariate normality. WIREs Computational Statistics, 6(1):56–73, January/February 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Crutchfield:2014:DT

[292] James P. Crutchfield. The dreams of theory. WIREs Computational Statistics, 6(2):75–79, March/April 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Leonard:2014:PHB

[293] Thomas Hoskyns Leonard. A personal history of Bayesian statistics. WIREs Computational Statistics, 6(2):80–115, March/April 2014. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2014:LAR

[294] Hongyang Zhang and Ruben H. Zamar. Least angle regression for model selection. WIREs Computational Statistics, 6(2):116–123, March/April 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2014:RMB

[295] Waldyn Martinez and J. Brian Gray. The role of margins in boosting and ensemble performance. WIREs Computational Statistics, 6(2):124– 131, March/April 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Iverson:2014:GFI

[296] Todd Iverson. Generalized fiducial inference. WIREs Computational Statistics, 6(2):132–143, March/April 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Coppersmith:2014:VN

[297] Glen Coppersmith. Vertex nomination. WIREs Computational Statistics, 6(2):144–153, March/April 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Schiff:2014:SMC

[298] Joel L. Schiff. Scientific modeling with cellular automata. WIREs Computational Statistics, 6(3):155–163, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Virius:2014:OOC

[299] Miroslav Virius and Rudolf Pecinovský. Object oriented computing. WIREs Computational Statistics, 6(3):164–176, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Qiao:2014:VSU

[300] Xingye Qiao. Variable selection using L_{q} penalties. WIREs Computational Statistics, 6(3):177–184, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kosmidis:2014:BPE

[301] Ioannis Kosmidis. Bias in parametric estimation: reduction and useful side-effects. WIREs Computational Statistics, 6(3):185–196, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2014:VSV

[302] Lingsong Zhang and Yao Wang. Visualizing singular value decomposition. WIREs Computational Statistics, 6(3):197–201, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Phillips:2014:T

[303] Daisy Phillips. Tessellation. WIREs Computational Statistics, 6(3):202– 209, May/June 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rossman:2014:USB

[304] Allan J. Rossman and Beth L. Chance. Using simulation-based inference for learning introductory statistics. WIREs Computational Statistics, 6 (4):211–221, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hoerl:2014:AST

[305] Roger W. Hoerl, Ronald D. Snee, and Richard D. De Veaux. Applying statistical thinking to ‘Big Data’ problems. WIREs Computational Statistics, 6(4):222–232, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Levine:2014:JCG

[306] Richard A. Levine, Eric Sampson, and Thomas C. M. Lee. Journal of Computational and Graphical Statistics. WIREs Computational Statistics, 6(4):233–239, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liang:2014:OSA

[307] Faming Liang. An overview of stochastic approximation Monte Carlo. WIREs Computational Statistics, 6(4):240–254, July/August 2014. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tong:2014:EVC

[308] Tiejun Tong, Cheng Wang, and Yuedong Wang. Estimation of variances and covariances for high-dimensional data: a selective review. WIREs Computational Statistics, 6(4):255–264, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Efromovich:2014:NRM

[309] Sam Efromovich. Nonparametric regression with missing data. WIREs Computational Statistics, 6(4):265–275, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lee:2014:BSE

[310] Sik-Yum Lee and Xin-Yuan Song. Bayesian structural equation model. WIREs Computational Statistics, 6(4):276–287, July/August 2014. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hu:2014:ARB

[311] Feifang Hu, Yanqing Hu, Zhenjun Ma, and William F. Rosenberger. Adaptive randomization for balancing over covariates. WIREs Computational Statistics, 6(4):288–303, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chow:2014:BBD

[312] Shein-Chung Chow. Bioavailability and bioequivalence in drug development. WIREs Computational Statistics, 6(4):304–312, July/August 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Winkler:2014:MRL

[313] William E. Winkler. Matching and record linkage. WIREs Computational Statistics, 6(5):313–325, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ensor:2014:CFC

[314] Katherine Bennett Ensor and Ginger M. Koev. Computational finance: correlation, volatility, and markets. WIREs Computational Statistics, 6 (5):326–340, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Duffy:2014:WDC

[315] Austen C. Duffy. Where do computational mathematics and computational statistics converge? WIREs Computational Statistics, 6(5):341– 351, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gilli:2014:OC

[316] Manfred Gilli and Enrico Schumann. Optimization cultures. WIREs Computational Statistics, 6(5):352–358, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shyr:2014:ADA

[317] Jing Shyr and Damir Spisic. Automated data analysis. WIREs Computational Statistics, 6(5):359–366, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Goodman:2014:ESI

[318] Arnold Goodman. Evolution of symposia on the interface of computing and statistics defines data science to be the interface. WIREs Computational Statistics, 6(5):367–377, September/October 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Thompson:2014:SBE

[319] James R. Thompson. Simulation-based estimation: a case study in oncology (SIMEST) and a case study in portfolio selection (SIMU-GRAM). WIREs Computational Statistics, 6(6):379–385, November/ December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kroese:2014:WMC

[320] Dirk P. Kroese, Tim Brereton, Thomas Taimre, and Zdravko I. Botev. Why the Monte Carlo method is so important today. WIREs Computational Statistics, 6(6):386–392, November/December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2014:VAH

[321] Yazhen Wang and Jian Zou. Volatility analysis in high-frequency financial data. WIREs Computational Statistics, 6(6):393–404, November/ December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lawson:2014:HMS

[322] Andrew B. Lawson. Hierarchical modeling in spatial epidemiology. WIREs Computational Statistics, 6(6):405–417, November/December 2014. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dannemann:2014:SHM

[323] Jörn Dannemann, Hajo Holzmann, and Anna Leister. Semiparametric hidden Markov models: identifiability and estimation. WIREs Computational Statistics, 6(6):418–425, November/December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Harenberg:2014:CDL

[324] Steve Harenberg, Gonzalo Bello, L. Gjeltema, Stephen Ranshous, Jitendra Harlalka, Ramona Seay, Kanchana Padmanabhan, and Nagiza Samatova. Community detection in large-scale networks: a survey and empirical evaluation. WIREs Computational Statistics, 6(6):426–439, November/December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tondel:2014:ACM

[325] Kristin Tøndel and Harald Martens. Analyzing complex mathematical model behavior by partial least squares regression-based multivariate metamodeling. WIREs Computational Statistics, 6(6):440–475, November/December 2014. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gijbels:2015:VSU

[326] Irène Gijbels, Anneleen Verhasselt, and Inge Vrinssen. Variable selection using P-splines. WIREs Computational Statistics, 7(1):1–20, January/ February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2015:BMF

[327] Linlin Zhang, Michele Guindani, and Marina Vannucci. Bayesian models for functional magnetic resonance imaging data analysis. WIREs Computational Statistics, 7(1):21–41, January/February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gower:2015:BQD

[328] John C. Gower, Niël J. Le Roux, and Sugnet Gardner-Lubbe. Biplots: quantitative data. WIREs Computational Statistics, 7(1):42–62, January/ February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chenouri:2015:RDR

[329] Shojaeddin Chenouri, Jiaxi Liang, and Christopher G. Small. Robust dimension reduction. WIREs Computational Statistics, 7(1):63–69, January/February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ma:2015:LBD

[330] Ping Ma and Xiaoxiao Sun. Leveraging for big data regression. WIREs Computational Statistics, 7(1):70–76, January/February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liu:2015:FPF

[331] Chuanhai Liu and Ryan Martin. Frameworks for prior-free posterior probabilistic inference. WIREs Computational Statistics, 7(1):77–85, January/ February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wikle:2015:MPS

[332] Christopher K. Wikle. Modern perspectives on statistics for spatiotemporal data. WIREs Computational Statistics, 7(1):86–98, January/ February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhao:2015:IAO

[333] Qing Zhao, Xingjie Shi, Jian Huang, Jin Liu, Yang Li, and Shuangge Ma. Integrative analysis of ‘-omics’ data using penalty functions. WIREs Computational Statistics, 7(1):99–108, January/February 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Marron:2015:DWD

[334] J. S. Marron. Distance-weighted discrimination. WIREs Computational Statistics, 7(2):109–114, March/April 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hu:2015:VLG

[335] Yifan Hu and Lei Shi. Visualizing large graphs. WIREs Computational Statistics, 7(2):115–136, March/April 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gromping:2015:VIR

[336] Ulrike Grömping. Variable importance in regression models. WIREs Computational Statistics, 7(2):137–152, March/April 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See comments [424].

Turlach:2015:QP

[337] Berwin A. Turlach and Stephen J. Wright. Quadratic programming. WIREs Computational Statistics, 8(4):153–159, March/April 2015. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bekkerman:2015:RSE

[338] Anton Bekkerman. The role of simulations in econometrics pedagogy. WIREs Computational Statistics, 8(4):160–165, March/April 2015. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Baumer:2015:RM

[339] Benjamin Baumer and Dana Udwin. R Markdown. WIREs Computational Statistics, 7(3):167–177, May/June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhong:2015:TSD

[340] Wenxuan Zhong, Xin Xing, and Kenneth Suslick. Tensor sufficient dimension reduction. WIREs Computational Statistics, 7(3):178–184, May/June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ghosh:2015:BMS

[341] Joyee Ghosh. Bayesian model selection using the median probability model. WIREs Computational Statistics, 7(3):185–193, May/June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shults:2015:QLS

[342] Justine Shults. Quasi-least squares fitting. WIREs Computational Statistics, 7(3):194–204, May/June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Durre:2015:REP

[343] Alexander Dürre, Roland Fried, and Tobias Liboschik. Robust estimation of (partial) autocorrelation. WIREs Computational Statistics, 7(3):205– 222, May/June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ranshous:2015:ADD

[344] Stephen Ranshous, Shitian Shen, Danai Koutra, Steve Harenberg, Christos Faloutsos, and Nagiza F. Samatova. Anomaly detection in dynamic networks: a survey. WIREs Computational Statistics, 7(3):223–247, May/ June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ding:2015:HOS

[345] Shanshan Ding and R. Dennis Cook. Higher-order sliced inverse regressions. WIREs Computational Statistics, 7(4):249–257, July/August 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martinez:2015:ICL

[346] Angel Martinez and Wendy Martinez. At the interface of computational linguistics and statistics. WIREs Computational Statistics, 7(4):258–274, July/August 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Regier:2015:SUF

[347] Michael David Regier and R. David Parker. Smoothing using fractional polynomials: an alternative to polynomials and splines in applied research. WIREs Computational Statistics, 7(4):275–283, July/August 2015. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bianchi:2015:ABM

[348] Federico Bianchi and Flaminio Squazzoni. Agent-based models in sociology. WIREs Computational Statistics, 7(4):284–306, July/August 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gelfand:2015:BWF

[349] Alan E. Gelfand and Sudipto Banerjee. Bayesian wombling: finding rapid change in spatial maps. WIREs Computational Statistics, 7(5):307–315, September/October 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Albert:2015:PEU

[350] Jim Albert. Player evaluation using win probabilities in sports competitions. WIREs Computational Statistics, 7(5):316–325, September/ October 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Karl:2015:PGT

[351] Andrew Karl, James Wisnowski, and W. Heath Rushing. A practical guide to text mining with topic extraction. WIREs Computational Statistics, 7(5):326–340, September/October 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Qiao:2015:LOD

[352] Xingye Qiao. Learning ordinal data. WIREs Computational Statistics, 7(5):341–346, September/October 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Domeniconi:2015:BCC

[353] Carlotta Domeniconi and Kathryn Laskey. Bayesian co-clustering. WIREs Computational Statistics, 7(5):347–356, September/October 2015. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhu:2015:UMV

[354] Mu Zhu. Use of majority votes in statistical learning. WIREs Computational Statistics, 7(6):357–371, November/December 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Medina:2015:RSS

[355] Marco Avella Medina and Elvezio Ronchetti. Robust statistics: a selective overview and new directions. WIREs Computational Statistics, 7(6):372– 393, November/December 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shen:2015:LM

[356] Dan Shen, D. P. Dittmer, and J. S. Marron. The limits of multiplexing. WIREs Computational Statistics, 7(6):394–399, November/December 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Suwan:2015:BVN

[357] Shakira Suwan, Dominic S. Lee, and Carey E. Priebe. Bayesian vertex nomination using content and context. WIREs Computational Statistics, 7(6):400–416, November/December 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IIa

[358] Anonymous. Issue information. WIREs Computational Statistics, 8(1): 1–3, January/February 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tendick:2016:SMC

[359] Patrick H. Tendick, Lorraine Denby, and Wen-Hua Ju. Statistical methods for complex event processing and real time decision making. WIREs Computational Statistics, 8(1):5–26, January/February 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rizzo:2016:ED

[360] Maria L. Rizzo and Gábor J. Székely. Energy distance. WIREs Computational Statistics, 8(1):27–38, January/February 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kott:2016:CWS

[361] Phillip S. Kott. Calibration weighting in survey sampling. WIREs Computational Statistics, 8(1):39–53, January/February 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sun:2016:LRI

[362] Yan Sun. Linear regression with interval-valued data. WIREs Computational Statistics, 8(1):54–60, January/February 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tong:2016:CIV

[363] Xin Tong, Yang Feng, and Anqi Zhao. Cover image, volume 8, issue 2. WIREs Computational Statistics, 8(2):i, March/April 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IIb

[364] Anonymous. Issue information. WIREs Computational Statistics, 8(2): 61–63, March/April 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tong:2016:SNP

[365] Xin Tong, Yang Feng, and Anqi Zhao. A survey on Neyman–Pearson classification and suggestions for future research. WIREs Computational Statistics, 8(2):64–81, March/April 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gower:2016:BQD

[366] John C. Gower, Niël J. Le Roux, and Sugnet Gardner-Lubbe. Biplots: qualititative data. WIREs Computational Statistics, 8(2):82–111, March/ April 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hu:2016:SMS

[367] Xiangping Hu and Ingelin Steinsland. Spatial modeling with system of stochastic partial differential equations. WIREs Computational Statistics, 8(2):112–125, March/April 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IIc

[368] Anonymous. Issue information. WIREs Computational Statistics, 8(3): 127–129, May/June 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dudek:2016:PAM

[369] Anna E. Dudek, Harry Hurd, and Wioletta Wójtowicz. Periodic autoregressive moving average methods based on Fourier representation of periodic coefficients. WIREs Computational Statistics, 8(3):130–149, May/ June 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IId

[370] Anonymous. Issue information. WIREs Computational Statistics, 8(4): 151–153, July/August 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:VIR

[371] Anonymous. Variable importance in regression models. WIREs Computational Statistics, 8(4):154–157, July/August 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IIe

[372] Anonymous. Issue information. WIREs Computational Statistics, 8(5): 159–161, September/October 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Datta:2016:NNG

[373] Abhirup Datta, Sudipto Banerjee, Andrew O. Finley, and Alan E. Gelfand. On nearest-neighbor Gaussian process models for massive spatial data. WIREs Computational Statistics, 8(5):162–171, September/October 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Diday:2016:TCD

[374] Edwin Diday. Thinking by classes in data science: the symbolic data analysis paradigm. WIREs Computational Statistics, 8(5):172–205, September/October 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sampath:2016:TTP

[375] Srinath Sampath, Adriano Caloiaro, Wayne Johnson, and Joseph S. Verducci. The top-K tau-path screen for monotone association in subpopulations. WIREs Computational Statistics, 8(5):206–218, September/ October 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2016:IIf

[376] Anonymous. Issue information. WIREs Computational Statistics, 8(6): 219–221, November/December 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McDonald:2016:ASS

[377] Gary C. McDonald. Applications of subset selection procedures and Bayesian ranking methods in analysis of traffic fatality data. WIREs Computational Statistics, 8(6):222–237, November/December 2016. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IIa

[378] Anonymous. Issue information. WIREs Computational Statistics, 9(1): e1386:1–e1386:??, January/February 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wilson:2017:DRT

[379] Seunghye J. Wilson. Data representation for time series data mining: time domain approaches. WIREs Computational Statistics, 9(1):e1392:1– e1392:??, January/February 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IIb

[380] Anonymous. Issue information. WIREs Computational Statistics, 9(2): e1387:1–e1387:??, March/April 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Gentle:2017:EN

[381] James E. Gentle and David W. Scott. Editor’s note. WIREs Computational Statistics, 9(2):e1395:1–e1395:??, March/April 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Morgan:2017:OED

[382] J. P. Morgan and Jonathan Stallings. Optimal experimental design that targets meaningful information. WIREs Computational Statistics, 9(2): e1393:1–e1393:??, March/April 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Maity:2017:NFC

[383] Arnab Maity. Nonparametric functional concurrent regression models. WIREs Computational Statistics, 9(2):e1394:1–e1394:??, March/April 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IIc

[384] Anonymous. Issue information. WIREs Computational Statistics, 9(3): e1388:1–e1388:??, May/June 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Degras:2017:SCB

[385] David Degras. Simultaneous confidence bands for the mean of functional data. WIREs Computational Statistics, 9(3):e1397:1–e1397:??, May/June 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Inouye:2017:RMD

[386] David I. Inouye, Eunho Yang, Genevera I. Allen, and Pradeep Ravikumar. A review of multivariate distributions for count data derived from the Poisson distribution. WIREs Computational Statistics, 9(3):e1398:1–e1398:??, May/June 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Herranz:2017:URT

[387] Edward Herranz. Unit root tests. WIREs Computational Statistics, 9(3): e1396:1–e1396:??, May/June 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IId

[388] Anonymous. Issue information. WIREs Computational Statistics, 9(4): e1389:1–e1389:??, July/August 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Song:2017:SAF

[389] Kai-Sheng Song and Robert Kieschnick. Statistical analysis of financial data with many zeros. WIREs Computational Statistics, 9(4):e1399:1– e1399:??, July/August 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hung:2017:RGS

[390] Hung Hung and Henry Horng-Shing Lu. A review on the generalization of sufficient dimension reduction methods with the additional information. WIREs Computational Statistics, 9(4):e1401:1–e1401:??, July/ August 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhou:2017:CKM

[391] Mai Zhou and Yifan Yang. Constrained Kaplan–Meier curve and empirical likelihood. WIREs Computational Statistics, 9(4):e1400:1–e1400:??, July/August 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IIe

[392] Anonymous. Issue information. WIREs Computational Statistics, 9(5): e1390:1–e1390:??, September/October 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhao:2017:VSP

[393] Yize Zhao and Qi Long. Variable selection in the presence of missing data: imputation-based methods. WIREs Computational Statistics, 9(5): e1402:1–e1402:??, September/October 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhao:2017:STA

[394] Yunpeng Zhao. A survey on theoretical advances of community detection in networks. WIREs Computational Statistics, 9(5):e1403:1–e1403:??, September/October 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Deng:2017:DPR

[395] Lih-Yuan Deng and Dale Bowman. Developments in pseudo-random number generators. WIREs Computational Statistics, 9(5):e1404:1–e1404:??, September/October 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Billard:2017:HCH

[396] L. Billard and Jaejik Kim. Hierarchical clustering for histogram data. WIREs Computational Statistics, 9(5):e1405:1–e1405:??, September/October 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2017:IIf

[397] Anonymous. Issue information. WIREs Computational Statistics, 9(6): e1391:1–e1391:??, November/December 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chan:2017:NEE

[398] Joshua C. C. Chan, Daniel J. Henderson, Christopher F. Parmeter, and Justin L. Tobias. Nonparametric estimation in economics: Bayesian and frequentist approaches. WIREs Computational Statistics, 9(6):e1406:1– e1406:??, November/December 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chow:2017:ASA

[399] Shein-Chung Chow and Fuyu Song. Analytical similarity assessment. WIREs Computational Statistics, 9(6):e1407:1–e1407:??, November/December 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

ElRafey:2017:RAS

[400] Amr ElRafey and Janusz Wojtusiak. Recent advances in scaling-down sampling methods in machine learning. WIREs Computational Statistics, 9(6):e1414:1–e1414:??, November/December 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kuismin:2017:ECP

[401] Markku O. Kuismin and Mikko J. Sillanpää. Estimation of covariance and precision matrix, network structure, and a view toward systems biology. WIREs Computational Statistics, 9(6):e1415:1–e1415:??, November/ December 2017. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2018:IIa

[402] Anonymous. Issue information. WIREs Computational Statistics, 10(1): e1408:1–e1408:??, January/February 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2018:MCK

[403] Su Chen and Michael Pokojovy. Modern and classical k-sample omnibus tests. WIREs Computational Statistics, 10(1):e1418:1–e1418:??, January/ February 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zurbenko:2018:KZF

[404] Igor G. Zurbenko and Devin Smith. Kolmogorov–Zurbenko filters in spatiotemporal analysis. WIREs Computational Statistics, 10(1):e1419:1– e1419:??, January/February 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dey:2018:HTP

[405] Rajarshi Dey. Hypothesis tests with precedence probabilities and precedence-type tests. WIREs Computational Statistics, 10(1):e1417:1– e1417:??, January/February 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Song:2018:ORR

[406] Qifan Song. An overview of reciprocal L_{1}-regularization for high dimensional regression data. WIREs Computational Statistics, 10(1):e1416:1– e1416:??, January/February 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2018:IIb

[407] Anonymous. Issue information. WIREs Computational Statistics, 10(2): e1409:1–e1409:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Neath:2018:BMC

[408] Andrew A. Neath, Javier E. Flores, and Joseph E. Cavanaugh. Bayesian multiple comparisons and model selection. WIREs Computational Statistics, 10(2):e1420:1–e1420:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yuan:2018:SML

[409] Mengdie Yuan and Guoqing Diao. Sieve maximum likelihood estimation in generalized linear models with an unknown link function. WIREs Computational Statistics, 10(2):e1425:1–e1425:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sahinoglu:2018:CCR

[410] Mehmet Sahinoglu. CLOUD computing risk management for server-farm repair rates, consumer load cycle, server-farm repair crew count, and additional servers. WIREs Computational Statistics, 10(2):e1424:1–e1424:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liu:2018:KBM

[411] Ying Liu, Victor de la Pena, and Tian Zheng. Kernel-based measures of association. WIREs Computational Statistics, 10(2):e1422:1–e1422:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Crawford:2018:CMB

[412] Forrest W. Crawford, Lam Si Tung Ho, and Marc A. Suchard. Computational methods for birth-death processes. WIREs Computational Statistics, 10(2):e1423:1–e1423:??, March/April 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2018:IIc

[413] Anonymous. Issue information. WIREs Computational Statistics, 10(3): e1410:1–e1410:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Keller:2018:HPD

[414] Sallie Keller, Stephanie Shipp, Gizem Korkmaz, Emily Molfino, Joshua Goldstein, Vicki Lancaster, Bianica Pires, David Higdon, Daniel Chen, and Aaron Schroeder. Harnessing the power of data to support community-based research. WIREs Computational Statistics, 10(3): e1426:1–e1426:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hubert:2018:MCD

[415] Mia Hubert, Michiel Debruyne, and Peter J. Rousseeuw. Minimum covariance determinant and extensions. WIREs Computational Statistics, 10(3):e1421:1–e1421:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Carcea:2018:GAF

[416] Marcel D. Carcea. Gini autocovariance function used for time series with heavy-tail distributions. WIREs Computational Statistics, 10(3):e1428:1– e1428:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Calvetti:2018:IPR

[417] D. Calvetti and E. Somersalo. Inverse problems: From regularization to Bayesian inference. WIREs Computational Statistics, 10(3):e1427:1– e1427:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Berry:2018:PMP

[418] Kenneth J. Berry, Janis E. Johnston, Paul W. Mielke Jr., and Lindsay A. Johnston. Permutation methods. Part II. WIREs Computational Statistics, 10(3):e1429:1–e1429:??, May/June 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2018:IId

[419] Anonymous. Issue information. WIREs Computational Statistics, 10(4): e1411:1–e1411:??, July/August 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Stewart:2018:CSV

[420] Thomas G. Stewart, Donglin Zeng, and Michael C. Wu. Constructing support vector machines with missing data. WIREs Computational Statistics, 10(4):e1430:1–e1430:??, July/August 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Qin:2018:RQD

[421] Yingli Qin. A review of quadratic discriminant analysis for highdimensional data. WIREs Computational Statistics, 10(4):e1434:1– e1434:??, July/August 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dutilleul:2018:ETS

[422] Pierre Dutilleul. Estimation and testing for separable variance–covariance structures. WIREs Computational Statistics, 10(4):e1432:1–e1432:??, July/August 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2018:MRU

[423] Yen-Chi Chen. Modal regression using kernel density estimation: a review. WIREs Computational Statistics, 10(4):e1431:1–e1431:??, July/ August 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Thomas:2018:DPV

[424] D. Roland Thomas, Ernest Kwan, and Bruno D. Zumbo. In defense of Pratt’s variable importance axioms: a response to Gromping. WIREs Computational Statistics, 10(4):e1433:1–e1433:??, July/August 2018. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). See [336].

Anonymous:2018:IIe

[425] Anonymous. Issue information. WIREs Computational Statistics, 10(5): e1412:1–e1412:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Robert:2018:AMA

[426] Christian P. Robert, Víctor Elvira, Nick Tawn, and Changye Wu. Accelerating MCMC algorithms. WIREs Computational Statistics, 10(5): e1435:1–e1435:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Nordhausen:2018:ICA

[427] Klaus Nordhausen and Hannu Oja. Independent component analysis: a statistical perspective. WIREs Computational Statistics, 10(5):e1440:1– e1440:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Soyer:2018:KFS

[428] Refik Soyer. Kalman filtering and sequential Bayesian analysis. WIREs Computational Statistics, 10(5):e1438:1–e1438:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Klemela:2018:LST

[429] Jussi Klemelä. Level set tree methods. WIREs Computational Statistics, 10(5):e1436:1–e1436:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sweeney:2018:SCE

[430] James Sweeney, Michael Salter-Townshend, Tamsin Edwards, Caitlin E. Buck, and Andrew C. Parnell. Statistical challenges in estimating past climate changes. WIREs Computational Statistics, 10(5):e1437:1–e1437:??, September/October 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2018:IIf

[431] Anonymous. Issue information. WIREs Computational Statistics, 10(6): e1413:1–e1413:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Markatou:2018:WCV

[432] Marianthi Markatou, Georgios Afendras, and Claudio Agostinelli. Weighted cross validation in model selection. WIREs Computational Statistics, 10(6):e1439:1–e1439:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Holan:2018:TSC

[433] Scott H. Holan and Nalini Ravishanker. Time series clustering and classification via frequency domain methods. WIREs Computational Statistics, 10(6):e1444:1–e1444:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Perrakis:2018:BVS

[434] Konstantinos Perrakis and Ioannis Ntzoufras. Bayesian variable selection using the hyper-g prior in WinBUGS. WIREs Computational Statistics, 10(6):e1442:1–e1442:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bakka:2018:SMR

[435] Haakon Bakka, Håvard Rue, Geir-Arne Fuglstad, Andrea Riebler, David Bolin, Janine Illian, Elias Krainski, Daniel Simpson, and Finn Lindgren. Spatial modeling with R-INLA: a review. WIREs Computational Statistics, 10(6):e1443:1–e1443:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lawson:2018:BLM

[436] Andrew B. Lawson. Bayesian latent modeling of spatio-temporal variation in small-area health data. WIREs Computational Statistics, 10(6): e1441:1–e1441:??, November/December 2018. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IIa

[437] Anonymous. Issue information. WIREs Computational Statistics, 11(1): e1445:1–e1445:??, January/February 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huo:2019:AI

[438] Xiaoming Huo and Shanshan Cao. Aggregated inference. WIREs Computational Statistics, 11(1):e1451:1–e1451:??, January/February 2019. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2019:CNM

[439] Yong Wang. Computation of the nonparametric maximum likelihood estimate of a univariate log-concave density. WIREs Computational Statistics, 11(1):e1452:1–e1452:??, January/February 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zou:2019:CHD

[440] Hui Zou. Classification with high dimensional features. WIREs Computational Statistics, 11(1):e1453:1–e1453:??, January/February 2019. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IIb

[441] Anonymous. Issue information. WIREs Computational Statistics, 11(2): e1446:1–e1446:??, March/April 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Helwig:2019:SNM

[442] Nathaniel E. Helwig. Statistical nonparametric mapping: Multivariate permutation tests for location, correlation, and regression problems in neuroimaging. WIREs Computational Statistics, 11(2):e1457:1–e1457:??, March/April 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bar:2019:EBM

[443] Haim Bar and Kangyan Liu. Empirical Bayes methods in variable selection. WIREs Computational Statistics, 11(2):e1455:1–e1455:??, March/ April 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

He:2019:SOF

[444] Kevin He, Han Xu, and Jian Kang. A selective overview of feature screening methods with applications to neuroimaging data. WIREs Computational Statistics, 11(2):e1454:1–e1454:??, March/April 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IIc

[445] Anonymous. Issue information. WIREs Computational Statistics, 11(3): e1447:1–e1447:??, May/June 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cavanaugh:2019:AIC

[446] Joseph E. Cavanaugh and Andrew A. Neath. The Akaike information criterion: Background, derivation, properties, application, interpretation, and refinements. WIREs Computational Statistics, 11(3):e1460:1– e1460:??, May/June 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

vandeVelden:2019:DBC

[447] Michel van de Velden, Alfonso Iodice D’Enza, and Angelos Markos. Distance-based clustering of mixed data. WIREs Computational Statistics, 11(3):e1456:1–e1456:??, May/June 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Adachi:2019:FAL

[448] Kohei Adachi. Factor analysis: Latent variable, matrix decomposition, and constrained uniqueness formulations. WIREs Computational Statistics, 11(3):e1458:1–e1458:??, May/June 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IId

[449] Anonymous. Issue information. WIREs Computational Statistics, 11(4): e1448:1–e1448:??, July/August 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Song:2019:SDR

[450] Jun Song. On sufficient dimension reduction for functional data: Inverse moment-based methods. WIREs Computational Statistics, 11(4):e1459:1– e1459:??, July/August 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lyubchich:2019:IRA

[451] Vyacheslav Lyubchich, Nathaniel K. Newlands, Azar Ghahari, Tahir Mahdi, and Yulia R. Gel. Insurance risk assessment in the face of climate change: Integrating data science and statistics. WIREs Computational Statistics, 11(4):e1462:1–e1462:??, July/August 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Polson:2019:BRT

[452] Nicholas G. Polson and Vadim Sokolov. Bayesian regularization: From Tikhonov to horseshoe. WIREs Computational Statistics, 11(4):e1463:1– e1463:??, July/August 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2019:NDE

[453] Zhipeng Wang and David W. Scott. Nonparametric density estimation for high-dimensional data — algorithms and applications. WIREs Computational Statistics, 11(4):e1461:1–e1461:??, July/August 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IIe

[454] Anonymous. Issue information. WIREs Computational Statistics, 11(5): e1449:1–e1449:??, September/October 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Efromovich:2019:SAF

[455] Sam Efromovich. Statistical analysis of fMRI using wavelets: Big data, denoising, large-p-small-n matrices. WIREs Computational Statistics, 11 (5):e1467:1–e1467:??, September/October 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Martin:2019:CEP

[456] Donald E. K. Martin. Computation of exact probabilities associated with overlapping pattern occurrences. WIREs Computational Statistics, 11(5): e1477:1–e1477:??, September/October 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chi:2019:MCC

[457] Eric C. Chi and Tianxi Li. Matrix completion from a computational statistics perspective. WIREs Computational Statistics, 11(5):e1469:1– e1469:??, September/October 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Beh:2019:MMC

[458] Eric J. Beh and Rosaria Lombardo. Multiple and multiway correspondence analysis. WIREs Computational Statistics, 11(5):e1464:1–e1464:??, September/October 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2019:IIf

[459] Anonymous. Issue information. WIREs Computational Statistics, 11(6): e1450:1–e1450:??, November/December 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ghosh:2019:COS

[460] Joyee Ghosh. Cauchy and other shrinkage priors for logistic regression in the presence of separation. WIREs Computational Statistics, 11(6): e1478:1–e1478:??, November/December 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Reiser:2019:GFT

[461] Mark Reiser. Goodness-of-fit testing in sparse contingency tables when the number of variables is large. WIREs Computational Statistics, 11 (6):e1470:1–e1470:??, November/December 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sakamoto:2019:IVC

[462] Wataru Sakamoto. Inference on variance components near boundary in linear mixed effect models. WIREs Computational Statistics, 11(6): e1466:1–e1466:??, November/December 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Yu:2019:ARD

[463] Philip L. H. Yu, Jiaqi Gu, and Hang Xu. Analysis of ranking data. WIREs Computational Statistics, 11(6):e1483:1–e1483:??, November/December 2019. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IIa

[464] Anonymous. Issue information. WIREs Computational Statistics, 12(1): e1471:1–e1471:??, January/February 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ferraro:2020:SC

[465] Maria Brigida Ferraro and Paolo Giordani. Soft clustering. WIREs Computational Statistics, 12(1):e1480:1–e1480:??, January/February 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Johansen:2020:RAH

[466] Thomas Haugland Johansen, Kajsa Møllersen, Samuel Ortega, Himar Fabelo, Aday Garcia, Gustavo M. Callico, and Fred Godtliebsen. Recent advances in hyperspectral imaging for melanoma detection. WIREs Computational Statistics, 12(1):e1465:1–e1465:??, January/February 2020. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Taylor:2020:HDC

[467] Jerome Taylor and Kshitij Khare. High-dimensional covariance estimation for Gaussian directed acyclic graph models with given order. WIREs Computational Statistics, 12(1):e1468:1–e1468:??, January/ February 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ziebarth:2020:MGP

[468] Jesse D. Ziebarth and Yan Cui. Multiscalar genetic pathway modeling with hybrid Bayesian networks. WIREs Computational Statistics, 12(1): e1479:1–e1479:??, January/February 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IIb

[469] Anonymous. Issue information. WIREs Computational Statistics, 12(2): e1472:1–e1472:??, March/April 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cook:2020:EM

[470] R. Dennis Cook. Envelope methods. WIREs Computational Statistics, 12 (2):e1484:1–e1484:??, March/April 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Famoye:2020:RUB

[471] Felix Famoye and Carl Lee. Review of univariate and bivariate exponentiated exponential-geometric distributions. WIREs Computational Statistics, 12(2):e1481:1–e1481:??, March/April 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cheng:2020:TDD

[472] Yu-Hsiang Cheng, Tzee-Ming Huang, and Su-Yun Huang. Tensor decomposition for dimension reduction. WIREs Computational Statistics, 12 (2):e1482:1–e1482:??, March/April 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lam:2020:HDC

[473] Clifford Lam. High-dimensional covariance matrix estimation. WIREs Computational Statistics, 12(2):e1485:1–e1485:??, March/April 2020. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IIc

[474] Anonymous. Issue information. WIREs Computational Statistics, 12(3): e1473:1–e1473:??, May/June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Barbeito:2020:NCE

[475] Inés Barbeito and Ricardo Cao. Nonparametric curve estimation and bootstrap bandwidth selection. WIREs Computational Statistics, 12(3): e1488:1–e1488:??, May/June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cizek:2020:RNR

[476] Pavel Cízek and Serhan Sadikoglu. Robust nonparametric regression: a review. WIREs Computational Statistics, 12(3):e1492:1–e1492:??, May/ June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Li:2020:NCI

[477] Chun Li and Xiaodan Fan. On nonparametric conditional independence tests for continuous variables. WIREs Computational Statistics, 12(3): e1489:1–e1489:??, May/June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Klemela:2020:NVP

[478] Jussi Klemelä. Nonparametric volatility prediction. WIREs Computational Statistics, 12(3):e1491:1–e1491:??, May/June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Bharath:2020:ASD

[479] Karthik Bharath and Sebastian Kurtek. Analysis of shape data: From landmarks to elastic curves. WIREs Computational Statistics, 12(3): e1495:1–e1495:??, May/June 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IId

[480] Anonymous. Issue information. WIREs Computational Statistics, 12(4): e1474:1–e1474:??, July/August 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Vats:2020:AMC

[481] Dootika Vats, Nathan Robertson, James M. Flegal, and Galin L. Jones. Analyzing Markov chain Monte Carlo output. WIREs Computational Statistics, 12(4):e1501:1–e1501:??, July/August 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Grazian:2020:RAB

[482] Clara Grazian and Yanan Fan. A review of approximate Bayesian computation methods via density estimation: Inference for simulator-models. WIREs Computational Statistics, 12(4):e1486:1–e1486:??, July/August 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Golyandina:2020:PCS

[483] Nina Golyandina. Particularities and commonalities of singular spectrum analysis as a method of time series analysis and signal processing. WIREs Computational Statistics, 12(4):e1487:1–e1487:??, July/August 2020. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jiang:2020:FBA

[484] Ci-Ren Jiang and Lu-Hung Chen. Filtering-based approaches for functional data classification. WIREs Computational Statistics, 12(4):e1490:1– e1490:??, July/August 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IIe

[485] Anonymous. Issue information. WIREs Computational Statistics, 12(5): e1475:1–e1475:??, September–October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Fouedjio:2020:CMG

[486] Francky Fouedjio. Clustering of multivariate geostatistical data. WIREs Computational Statistics, 12(5):e1510:1–e1510:??, September–October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Xi:2020:AML

[487] Bowei Xi. Adversarial machine learning for cybersecurity and computer vision: Current developments and challenges. WIREs Computational Statistics, 12(5):e1511:1–e1511:??, September–October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Polson:2020:DLC

[488] Nicholas Polson and Vadim Sokolov. Deep learning: Computational aspects. WIREs Computational Statistics, 12(5):e1500:1–e1500:??, September–October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cai:2020:IIT

[489] Chencheng Cai, Rong Chen, and Min ge Xie. Individualized inference through fusion learning. WIREs Computational Statistics, 12(5):e1498:1– e1498:??, September–October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kim:2020:MMG

[490] Inyoung Kim, Liang Shan, Jiali Lin, Wenyu Gao, Byung-Jun Kim, and Hamdy Mahmoud. Multiple and multilevel graphical models. WIREs Computational Statistics, 12(5):e1497:1–e1497:??, September– October 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2020:IIf

[491] Anonymous. Issue information. WIREs Computational Statistics, 12(6): e1476:1–e1476:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rudra:2020:MTA

[492] Pratyaydipta Rudra, Efrén Cruz-Cortés, Xuhong Zhang, and Debashis Ghosh. Multiple testing approaches for hypotheses in integrative genomics. WIREs Computational Statistics, 12(6):e1493:1–e1493:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Washington:2020:AVA

[493] Benjamin J. Washington and Lynne Seymour. An adapted vector autoregressive expectation maximization imputation algorithm for climate data networks. WIREs Computational Statistics, 12(6):e1494:1–e1494:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Govindarajulu:2020:RCA

[494] Usha S. Govindarajulu and Ralph B. D’Agostino Sr. Review of current advances in survival analysis and frailty models. WIREs Computational Statistics, 12(6):e1504:1–e1504:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scharf:2020:AMM

[495] Henry R. Scharf and Frances E. Buderman. Animal movement models for multiple individuals. WIREs Computational Statistics, 12(6):e1506:1– e1506:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Blake:2020:NCE

[496] Tayler A. Blake and Yoonkyung Lee. Nonparametric covariance estimation with shrinkage toward stationary models. WIREs Computational Statistics, 12(6):e1507:1–e1507:??, November–December 2020. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IIa

[497] Anonymous. Issue information. WIREs Computational Statistics, 13(1): e1517:1–e1517:??, January–February 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Weiss:2021:SCT

[498] Christian H. Weiß. Stationary count time series models. WIREs Computational Statistics, 13(1):e1502:1–e1502:??, January–February 2021. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hurley:2021:MEU

[499] Catherine B. Hurley. Model exploration using conditional visualization. WIREs Computational Statistics, 13(1):e1503:1–e1503:??, January– February 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kokonendji:2021:TRM

[500] Célestin C. Kokonendji, Wagner H. Bonat, and Rahma Abid. Tweedie regression models and its geometric sums for (semi-)continuous data. WIREs Computational Statistics, 13(1):e1496:1–e1496:??, January– February 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cannings:2021:RPD

[501] Timothy I. Cannings. Random projections: Data perturbation for classification problems. WIREs Computational Statistics, 13(1):e1499:1– e1499:??, January–February 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Caudle:2021:RFF

[502] Kyle Caudle, Patrick S. Fleming, and Randy C. Hoover. A review of flow field forecasting: A high-dimensional forecasting procedure. WIREs Computational Statistics, 13(1):e1505:1–e1505:??, January–February 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shojaie:2021:CIV

[503] Ali Shojaie. Cover image, volume 13, issue 2. WIREs Computational Statistics, 13(2):e1552:1–e1552:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IIb

[504] Anonymous. Issue information. WIREs Computational Statistics, 13(2): e1518:1–e1518:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Eddelbuettel:2021:PCR

[505] Dirk Eddelbuettel. Parallel computing with R: A brief review. WIREs Computational Statistics, 13(2):e1515:1–e1515:??, March–April 2021. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ferreira:2021:BSS

[506] Marco A. R. Ferreira. Bayesian spatial and spatiotemporal models based on multiscale factorizations. WIREs Computational Statistics, 13(2): e1509:1–e1509:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shojaie:2021:DNA

[507] Ali Shojaie. Differential network analysis: a statistical perspective. WIREs Computational Statistics, 13(2):e1508:1–e1508:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Porcu:2021:YST

[508] Emilio Porcu, Reinhard Furrer, and Douglas Nychka. 30 years of space-time covariance functions. WIREs Computational Statistics, 13(2): e1512:1–e1512:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

McAndrew:2021:APE

[509] Thomas McAndrew, Nutcha Wattanachit, Graham C. Gibson, and Nicholas G. Reich. Aggregating predictions from experts: A review of statistical methods, experiments, and applications. WIREs Computational Statistics, 13(2):e1514:1–e1514:??, March–April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IIc

[510] Anonymous. Issue information. WIREs Computational Statistics, 13(3): e1519:1–e1519:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Scott:2021:E

[511] David W. Scott and James E. Gentle. Editorial. WIREs Computational Statistics, 13(3):e1555:1–e1555:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ravishanker:2021:IPH

[512] Nalini Ravishanker and Renjie Chen. An introduction to persistent homology for time series. WIREs Computational Statistics, 13(3):e1548:1– e1548:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2021:RSF

[513] Ying Chen, Thorsten Koch, Kian Guan Lim, Xiaofei Xu, and Nazgul Zakiyeva. A review study of functional autoregressive models with application to energy forecasting. WIREs Computational Statistics, 13(3): e1525:1–e1525:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huckemann:2021:DAN

[514] Stephan F. Huckemann and Benjamin Eltzner. Data analysis on nonstandard spaces. WIREs Computational Statistics, 13(3):e1526:1–e1526:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hang:2021:ASD

[515] Weiqiang Hang and Yingcun Xia. Advance of the sufficient dimension reduction. WIREs Computational Statistics, 13(3):e1516:1–e1516:??, May– June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dobelman:2021:SJR

[516] John A. Dobelman. Sampling James R. Thompson’s inspired nonparametric portfolio approaches. WIREs Computational Statistics, 13(3):e1542:1– e1542:??, May–June 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IId

[517] Anonymous. Issue information. WIREs Computational Statistics, 13(4): e1520:1–e1520:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lai:2021:RBG

[518] Wei-Ting Lai and Ray-Bing Chen. A review of Bayesian group selection approaches for linear regression models. WIREs Computational Statistics, 13(4):e1513:1–e1513:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Filzmoser:2021:RLR

[519] Peter Filzmoser and Klaus Nordhausen. Robust linear regression for highdimensional data: an overview. WIREs Computational Statistics, 13(4): e1524:1–e1524:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Suzuki:2021:WBR

[520] Joe Suzuki. Why BDeu? Regular Bayesian network structure learning with discrete and continuous variables. WIREs Computational Statistics, 13(4):e1554:1–e1554:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lin:2021:BMM

[521] Lin Lin and Boris P. Hejblum. Bayesian mixture models for cytometry data analysis. WIREs Computational Statistics, 13(4):e1535:1–e1535:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lysy:2021:CDB

[522] Martin Lysy, Masoud Asgharian, and Vahid Partovi Nia. A convergence diagnostic for Bayesian clustering. WIREs Computational Statistics, 13 (4):e1536:1–e1536:??, July–August 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IIe

[523] Anonymous. Issue information. WIREs Computational Statistics, 13(5): e1521:1–e1521:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Jin:2021:RLH

[524] Shaobo Jin and Youngjo Lee. A review of h-likelihood and hierarchical generalized linear model. WIREs Computational Statistics, 13(5): e1527:1–e1527:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Schmid:2021:CRA

[525] Matthias Schmid and Moritz Berger. Competing risks analysis for discrete time-to-event data. WIREs Computational Statistics, 13(5):e1529:1– e1529:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2021:ODT

[526] Haifeng Wang, Changzai Pan, Xiao Guo, Chunlin Ji, and Ke Deng. From object detection to text detection and recognition: a brief evolution history of optical character recognition. WIREs Computational Statistics, 13(5):e1547:1–e1547:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2021:MMC

[527] Jiaxin Zhang. Modern Monte Carlo methods for efficient uncertainty quantification and propagation: a survey. WIREs Computational Statistics, 13(5):e1539:1–e1539:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhou:2021:GWP

[528] Weiqiang Zhou and Hongkai Ji. Genome-wide prediction of chromatin accessibility based on gene expression. WIREs Computational Statistics, 13(5):e1544:1–e1544:??, September–October 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2021:IIf

[529] Anonymous. Issue information. WIREs Computational Statistics, 13(6): e1522:1–e1522:??, November–December 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kelter:2021:BFT

[530] Riko Kelter. Bayesian and frequentist testing for differences between two groups with parametric and nonparametric two-sample tests. WIREs Computational Statistics, 13(6):e1523:1–e1523:??, November–December 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sellers:2021:CMP

[531] Kimberly F. Sellers and Bailey Premeaux. Conway–Maxwell–Poisson regression models for dispersed count data. WIREs Computational Statistics, 13(6):e1533:1–e1533:??, November–December 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ling:2021:SRF

[532] Nengxiang Ling and Philippe Vieu. On semiparametric regression in functional data analysis. WIREs Computational Statistics, 13(6):e1538:1– e1538:??, November–December 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dattner:2021:DED

[533] Itai Dattner. Differential equations in data analysis. WIREs Computational Statistics, 13(6):e1534:1–e1534:??, November–December 2021. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tayefi:2021:COB

[534] Maryam Tayefi, Phuong Ngo, Taridzo Chomutare, Hercules Dalianis, Elisa Salvi, Andrius Budrionis, and Fred Godtliebsen. Challenges and opportunities beyond structured data in analysis of electronic health records. WIREs Computational Statistics, 13(6):e1549:1–e1549:??, November– December 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huser:2022:CI

[535] Raphaël Huser and Jennifer L. Wadsworth. Cover image. WIREs Computational Statistics, 14(1):e1577:1–e1577:??, January–February 2022. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IIa

[536] Anonymous. Issue information. WIREs Computational Statistics, 14(1): e1560:1–e1560:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Banks:2022:ARA

[537] David Banks, Víctor Gallego, Roi Naveiro, and David Ríos Insua. Adversarial risk analysis: an overview. WIREs Computational Statistics, 14(1): e1530:1–e1530:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Meyer:2022:CTP

[538] Renate Meyer, Matthew C. Edwards, Patricio Maturana-Russel, and Nelson Christensen. Computational techniques for parameter estimation of gravitational wave signals. WIREs Computational Statistics, 14(1): e1532:1–e1532:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Huser:2022:ASM

[539] Raphaël Huser and Jennifer L. Wadsworth. Advances in statistical modeling of spatial extremes. WIREs Computational Statistics, 14(1):e1537:1– e1537:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Young:2022:ZIMa

[540] Derek S. Young, Eric S. Roemmele, and Peng Yeh. Zero-inflated modeling part I: Traditional zero-inflated count regression models, their applications, and computational tools. WIREs Computational Statistics, 14(1): e1541:1–e1541:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Johnvictor:2022:CRB

[541] Anita Christaline Johnvictor, Vaishali Durgamahanthi, Ramya Meghana Pariti Venkata, and Nishtha Jethi. Critical review of bio-inspired optimization techniques. WIREs Computational Statistics, 14(1):e1528:1–e1528:??, January–February 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IIb

[542] Anonymous. Issue information. WIREs Computational Statistics, 14(2): e1561:1–e1561:??, March–April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hori:2022:IRTa

[543] Kazuki Hori, Hirotaka Fukuhara, and Tsuyoshi Yamada. Item response theory and its applications in educational measurement. Part I: Item response theory and its implementation in R. WIREs Computational Statistics, 14(2):e1531:1–e1531:??, March–April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Dey:2022:CDC

[544] Asim K. Dey, Yahui Tian, and Yulia R. Gel. Community detection in complex networks: From statistical foundations to data science applications. WIREs Computational Statistics, 14(2):e1566:1–e1566:??, March– April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Young:2022:ZIMb

[545] Derek S. Young, Eric S. Roemmele, and Xuan Shi. Zero-inflated modeling part II: Zero-inflated models for complex data structures. WIREs Computational Statistics, 14(2):e1540:1–e1540:??, March–April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tutz:2022:ORR

[546] Gerhard Tutz. Ordinal regression: A review and a taxonomy of models. WIREs Computational Statistics, 14(2):e1545:1–e1545:??, March– April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Park:2022:IGS

[547] Taeyoung Park and Seunghan Lee. Improving the Gibbs sampler. WIREs Computational Statistics, 14(2):e1546:1–e1546:??, March–April 2022. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IIc

[548] Anonymous. Issue information. WIREs Computational Statistics, 14(3): e1562:1–e1562:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hori:2022:IRTb

[549] Kazuki Hori, Hirotaka Fukuhara, and Tsuyoshi Yamada. Item response theory and its applications in educational measurement. Part II: Theory and practices of test equating in item response theory. WIREs Computational Statistics, 14(3):e1543:1–e1543:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2022:ICM

[550] Xiaoyu Zhang, Zhenwei Zhou, Hanfei Xu, and Ching-Ti Liu. Integrative clustering methods for multi-omics data. WIREs Computational Statistics, 14(3):e1553:1–e1553:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Grosser:2022:COR

[551] Joshua Größer and Ostap Okhrin. Copulae: An overview and recent developments. WIREs Computational Statistics, 14(3):e1557:1–e1557:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Sato-Ilic:2022:CSP

[552] Mika Sato-Ilic. Cluster-scaled principal component analysis. WIREs Computational Statistics, 14(3):e1572:1–e1572:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Price:2022:DCM

[553] Bradley S. Price, Corban Allenbrand, and Ben Sherwood. Detecting clusters in multivariate response regression. WIREs Computational Statistics, 14(3):e1551:1–e1551:??, May–June 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IId

[554] Anonymous. Issue information. WIREs Computational Statistics, 14(4): e1563:1–e1563:??, July–August 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Pan:2022:RSO

[555] Yan Pan, Markus Matilainen, Sara Taskinen, and Klaus Nordhausen. A review of second-order blind identification methods. WIREs Computational Statistics, 14(4):e1550:1–e1550:??, July–August 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ellis:2022:SRS

[556] Dorothy Ellis, Dongyuan Wu, and Susmita Datta. SAREV: a review on statistical analytics of single-cell RNA sequencing data. WIREs Computational Statistics, 14(4):e1558:1–e1558:??, July–August 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Nolan:2022:CAS

[557] John P. Nolan. Computational aspects of stable distributions. WIREs Computational Statistics, 14(4):e1569:1–e1569:??, July–August 2022. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Lee:2022:SAT

[558] Stuart Lee, Dianne Cook, Natalia da Silva, Ursula Laa, Nicholas Spyrison, Earo Wang, and H. Sherry Zhang. The state-of-the-art on tours for dynamic visualization of high-dimensional data. WIREs Computational Statistics, 14(4):e1573:1–e1573:??, July–August 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wang:2022:RLD

[559] Zhenbang Wang, Emanuel Ben-David, Guoqing Diao, and Martin Slawski. Regression with linked datasets subject to linkage error. WIREs Computational Statistics, 14(4):e1570:1–e1570:??, July–August 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IIe

[560] Anonymous. Issue information. WIREs Computational Statistics, 14(5): e1564:1–e1564:??, September–October 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Golosnoy:2022:SRM

[561] Vasyl Golosnoy, Bastian Gribisch, and Miriam Isabel Seifert. Sample and realized minimum variance portfolios: Estimation, statistical inference, and tests. WIREs Computational Statistics, 14(5):e1556:1–e1556:??, September–October 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

So:2022:VDD

[562] Mike K. P. So, Amanda M. Y. Chu, Cliff C. Y. Lo, and Chun Yin Ip. Volatility and dynamic dependence modeling: Review, applications, and financial risk management. WIREs Computational Statistics, 14(5): e1567:1–e1567:??, September–October 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kim:2022:PIP

[563] Taeho Kim, Benjamin Lieberman, George Luta, and Edsel A. Peña. Prediction intervals for Poisson-based regression models. WIREs Computational Statistics, 14(5):e1568:1–e1568:??, September–October 2022. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chen:2022:PSO

[564] Ping-Yang Chen, Ray-Bing Chen, and Weng Kee Wong. Particle swarm optimization for searching efficient experimental designs: a review. WIREs Computational Statistics, 14(5):e1578:1–e1578:??, September– October 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Datta:2022:NNS

[565] Abhirup Datta. Nearest-neighbor sparse Cholesky matrices in spatial statistics. WIREs Computational Statistics, 14(5):e1574:1–e1574:??, September–October 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2022:IIf

[566] Anonymous. Issue information. WIREs Computational Statistics, 14(6): e1565:1–e1565:??, November–December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Nash:2022:FMN

[567] John C. Nash. Function minimization and nonlinear least squares in R. WIREs Computational Statistics, 14(6):e1580:1–e1580:??, November– December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Soyer:2022:BMM

[568] Refik Soyer and Di Zhang. Bayesian modeling of multivariate time series of counts. WIREs Computational Statistics, 14(6):e1559:1–e1559:??, November–December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Liu:2022:SEU

[569] Tianmou Liu, Han Yu, and Rachael Hageman Blair. Stability estimation for unsupervised clustering: a review. WIREs Computational Statistics, 14(6):e1575:1–e1575:??, November–December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Tsai:2022:JGG

[570] Katherine Tsai, Oluwasanmi Koyejo, and Mladen Kolar. Joint Gaussian graphical model estimation: a survey. WIREs Computational Statistics, 14(6):e1582:1–e1582:??, November–December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cozzolino:2022:DC

[571] Irene Cozzolino and Maria Brigida Ferraro. Document clustering. WIREs Computational Statistics, 14(6):e1588:1–e1588:??, November–December 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2023:IIa

[572] Anonymous. Issue information. WIREs Computational Statistics, 15(1): e1589:1–e1589:??, January–February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Linero:2023:HWB

[573] Antonio R. Linero and Joseph L. Antonelli. The how and why of Bayesian nonparametric causal inference. WIREs Computational Statistics, 15(1): e1583:1–e1583:??, January–February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Llorente:2023:SUP

[574] Fernando Llorente, Luca Martino, Ernesto Curbelo, Javier López-Santiago, and David Delgado. On the safe use of prior densities for Bayesian model selection. WIREs Computational Statistics, 15(1): e1595:1–e1595:??, January–February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Kurihara:2023:EAS

[575] Koji Kurihara and Fumio Ishioka. Echelon analysis and its software for spatial lattice data. WIREs Computational Statistics, 15(1):e1579:1– e1579:??, January–February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Shi:2023:DIC

[576] Xu Shi, Ziyang Pan, and Wang Miao. Data integration in causal inference. WIREs Computational Statistics, 15(1):e1581:1–e1581:??, January– February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Swift:2023:RND

[577] Dionne Swift, Kellen Cresswell, Robert Johnson, Spiro Stilianoudakis, and Xingtao Wei. A review of normalization and differential abundance methods for microbiome counts data. WIREs Computational Statistics, 15(1):e1586:1–e1586:??, January–February 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2023:IIb

[578] Anonymous. Issue information. WIREs Computational Statistics, 15(2): e1590:1–e1590:??, March–April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zheng:2023:RAA

[579] Wanwan Zheng and Mingzhe Jin. A review on authorship attribution in text mining. WIREs Computational Statistics, 15(2):e1584:1–e1584:??, March–April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Cramer:2023:OCL

[580] Erhard Cramer. Ordered and censored lifetime data in reliability: an illustrative review. WIREs Computational Statistics, 15(2):e1571:1–e1571:??, March–April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Craigmile:2023:SIS

[581] Peter Craigmile, Radu Herbei, Ge Liu, and Grant Schneider. Statistical inference for stochastic differential equations. WIREs Computational Statistics, 15(2):e1585:1–e1585:??, March–April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Zhang:2023:PBT

[582] Jingyi Zhang, Ping Ma, Wenxuan Zhong, and Cheng Meng. Projectionbased techniques for high-dimensional optimal transport problems. WIREs Computational Statistics, 15(2):e1587:1–e1587:??, March–April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Chambers:2023:RRU

[583] Raymond L. Chambers, Enrico Fabrizi, Maria Giovanna Ranalli, Nicola Salvati, and Suojin Wang. Robust regression using probabilistically linked data. WIREs Computational Statistics, 15(2):e1596:1–e1596:??, March– April 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Anonymous:2023:IIc

[584] Anonymous. Issue information. WIREs Computational Statistics, 15 (3):e1591:1–e1591:??, May–June 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1591.

Jaeger:2023:CAM

[585] Adam Jaeger and David Banks. Cluster analysis: a modern statistical review. WIREs Computational Statistics, 15(3):e1597:1–e1597:??, May– June 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1597.

Duan:2023:SMC

[586] Jin-Chuan Duan, Shuping Li, and Yaxian Xu. Sequential Monte Carlo optimization and statistical inference. WIREs Computational Statistics, 15(3):e1598:1–e1598:??, May–June 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1598.

Liu:2023:RRA

[587] Pangpang Liu and Yichuan Zhao. A review of recent advances in empirical likelihood. WIREs Computational Statistics, 15(3):e1599:1–e1599:??, May–June 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1599.

Murakami:2023:UBW

[588] Hidetoshi Murakami and Seong-Keon Lee. On unbiasedness and biasedness of the Wilcoxon and some nonparametric tests. WIREs Computational Statistics, 15(3):e1600:1–e1600:??, May–June 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1600.

Tomita:2023:DVS

[589] Makoto Tomita. Data validation and statistical issues such as power and other considerations in genome-wide association study (GWAS). WIREs Computational Statistics, 15(3):e1601:1–e1601:??, May–June 2023. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

//wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1601.

Anonymous:2023:IId

[590] Anonymous. Issue information. WIREs Computational Statistics, 15(4): e1592:1–e1592:??, July–August 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1592.

Ahmed:2023:SST

[591] Syed Ejaz Ahmed, Dursun Aydin, and Ersin Yilmaz. A survey of smoothing techniques based on a backfitting algorithm in estimation of semiparametric additive models. WIREs Computational Statistics, 15(4): e1605:1–e1605:??, July–August 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1605.

Zhao:2023:SNA

[592] Yujie Zhao and Xiaoming Huo. A survey of numerical algorithms that can solve the lasso problems. WIREs Computational Statistics, 15(4): e1602:1–e1602:??, July–August 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1602.

Miecznikowski:2023:ECT

[593] Jeffrey C. Miecznikowski and Jiefei Wang. Error control in tree structured hypothesis testing. WIREs Computational Statistics, 15(4):e1603:1– e1603:??, July–August 2023. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1603.

Kubota:2023:DMS

[594] Takafumi Kubota. Diseases maps of spatial epidemiological data by R. WIREs Computational Statistics, 15(4):e1604:1–e1604:??, July– August 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1604.

Anonymous:2023:IIe

[595] Anonymous. Issue information. WIREs Computational Statistics, 15 (5):e1593:1–e1593:??, September–October 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1593.

Zhang:2023:ICM

[596] Jiawei Zhang, Yuhong Yang, and Jie Ding. Information criteria for model selection. WIREs Computational Statistics, 15(5):e1607:1– e1607:??, September–October 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1607.

Lazar:2023:FNE

[597] Nicole A. Lazar. Functional neuroimaging in the era of Big Data and Open Science: a modern overview. WIREs Computational Statistics, 15(5):e1609:1–e1609:??, September–October 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1609.

Conard:2023:SEI

[598] Ashley Mae Conard, Alan DenAdel, and Lorin Crawford. A spectrum of explainable and interpretable machine learning approaches for genomic studies. WIREs Computational Statistics, 15(5):e1617:1–e1617:??, September–October 2023. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1617.

Drake:2023:NSA

[599] Daniel F. Drake, Gordana Derado, Lijun Zhang, F. DuBois Bowman, and for the Alzheimer’s Disease Neuroimaging Initiative. Neuroimaging statistical approaches for determining neural correlates of Alzheimer’s disease via positron emission tomography imaging. WIREs Computational Statistics, 15(5):e1606:1–e1606:??, September–October 2023. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

//wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1606.

Anonymous:2023:IIf

[600] Anonymous. Issue information. WIREs Computational Statistics, 15 (6):e1594:1–e1594:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1594.

Lan:2023:SCC

[601] Shiwei Lan and Lulu Kang. Sampling constrained continuous probability distributions: a review. WIREs Computational Statistics, 15(6):e1608:1– e1608:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1608.

Wang:2023:TLM

[602] Hsiuying Wang. Tolerance limits for mixture-of-normal distributions with application to COVID-19 data. WIREs Computational Statistics, 15(6):e1611:1–e1611:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1611.

Maden:2023:RMS

[603] Engin Maden and Pinar Karagoz. Recent methods on short text stream clustering: a survey study. WIREs Computational Statistics, 15(6):e1610:1–e1610:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1610.

Wang:2023:SAA

[604] Jingshen Wang and Xuming He. Subgroup analysis and adaptive experiments crave for debiasing. WIREs Computational Statistics, 15(6):e1614:1–e1614:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1614.

Williams:2023:PLF

[605] Aaron R. Williams and Claire McKay Bowen. The promise and limitations of formal privacy. WIREs Computational Statistics, 15(6):e1615:1– e1615:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1615.

Baumer:2023:BIS

[606] Benjamin S. Baumer, Gregory J. Matthews, and Quang Nguyen. Big ideas in sports analytics and statistical tools for their investigation. WIREs Computational Statistics, 15(6):e1612:1–e1612:??, November– December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1612.

Kang:2023:CIP

[607] Joseph Kang, Darcy Steeg Morris, Patrick Joyce, and Isaac Dompreh. On calibrated inverse probability weighting and generalized boosting propensity score models for mean estimation with incomplete survey data. WIREs Computational Statistics, 15(6):e1616:1–e1616:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1616.

Franco:2023:CSS

[608] Carolina Franco and Poulami Maitra. Combining surveys in small area estimation using area-level models. WIREs Computational Statistics, 15(6):e1613:1–e1613:??, November–December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1613.

Kuroda:2023:AEA

[609] Masahiro Kuroda and Zhi Geng. Acceleration of the EM algorithm. WIREs Computational Statistics, 15(6):e1618:1–e1618:??, November– December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1618.

Madsen:2023:RMM

[610] Lisa Madsen and J. Andrew Royle. A review of N-mixture models. WIREs Computational Statistics, 15(6):e1625:1–e1625:??, November– December 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1625.

Anonymous:2024:IIa

[611] Anonymous. Issue information. WIREs Computational Statistics, 16(1): e1619:1–e1619:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1619.

Hornung:2024:PAP

[612] Roman Hornung, Frederik Ludwigs, Jonas Hagenberg, and Anne-Laure Boulesteix. Prediction approaches for partly missing multiomics covariate data: a literature review and an empirical comparison study. WIREs Computational Statistics, 16(1):e1626:1–e1626:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-

5108 (electronic). URL https://wires.onlinelibrary.wiley.com/
doi/epdf/10.1002/wics.1626.

Meyer:2024:AEI

[613] Mark J. Meyer. Advances in estimation and inference for historical functional linear models. WIREs Computational Statistics, 16(1):e1627:1– e1627:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1627.

Wang:2024:SCP

[614] Haoyun Wang and Yao Xie. Sequential change-point detection: Computation versus statistical performance. WIREs Computational Statistics, 16(1):e1628:1–e1628:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1628.

Megahed:2024:CRG

[615] Mohammed Megahed and Ammar Mohammed. A comprehensive review of generative adversarial networks: Fundamentals, applications, and challenges. WIREs Computational Statistics, 16(1):e1629:1–e1629:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-

5108 (electronic). URL https://wires.onlinelibrary.wiley.com/
doi/epdf/10.1002/wics.1629.

Govindasamy:2024:SRE

[616] Priyalatha Govindasamy, Nor Junainah Mohd Isa, Nor Firdous Mohamed, Amelia Mohd Noor, Lin Ma, Antonio Olmos, and Kathy Green. A systematic review of exploratory factor analysis packages in R software. WIREs Computational Statistics, 16(1):e1630:1–e1630:??, January– February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1630.

Stacy:2024:BTM

[617] Stephanie Stacy, Siyi Gong, Aishni Parab, Minglu Zhao, Kaiwen Jiang, and Tao Gao. A Bayesian theory of mind approach to modeling cooperation and communication. WIREs Computational Statistics, 16(1): e1631:1–e1631:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1631.

Hunter:2024:UCU

[618] David R. Hunter. Unsupervised clustering using nonparametric finite mixture models. WIREs Computational Statistics, 16(1):e1632:1–e1632:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-

5108 (electronic). URL https://wires.onlinelibrary.wiley.com/
doi/epdf/10.1002/wics.1632.

Wu:2024:BEM

[619] Stephanie M. Wu and Briana Joy K. Stephenson. Bayesian estimation methods for survey data with potential applications to health disparities research. WIREs Computational Statistics, 16(1):e1633:1–e1633:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-

5108 (electronic). URL https://wires.onlinelibrary.wiley.com/
doi/epdf/10.1002/wics.1633.

Beutner:2024:DMA

[620] Eric Beutner. Delta method, asymptotic distribution. WIREs Computational Statistics, 16(1):e1634:1–e1634:??, January–February 2024. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

//wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1634.

Miao:2024:SMG

[621] Jiacheng Miao, Yixuan Wu, and Qiongshi Lu. Statistical methods for gene-environment interaction analysis. WIREs Computational Statistics, 16(1):e1635:1–e1635:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1635.

Hu:2024:AMP

[622] Jingchen Hu and Claire McKay Bowen. Advancing microdata privacy protection: a review of synthetic data methods. WIREs Computational Statistics, 16(1):e1636:1–e1636:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1636.

Hung:2024:RMC

[623] Ying-Chao Hung. A review of Monte Carlo and quasi-Monte Carlo sampling techniques. WIREs Computational Statistics, 16(2):e1637:1– e1637:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1637. See correction [637].

Wang:2024:RFD

[624] Shuoyang Wang, Yuan Huang, and Guanqun Cao. Review on functional data classification. WIREs Computational Statistics, 16(2):e1638:1– e1638:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1638.

Chu:2024:RCA

[625] Haitao Chu, Lifeng Lin, Zheng Wang, Zilin Wang, Yong Chen, and Joseph C. Cappelleri. A review and comparison of arm-based versus contrast-based network meta-analysis for binary outcomes-understanding their differences and limitations. WIREs Computational Statistics, 16(2): e1639:1–e1639:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1639.

Haghbin:2024:JUM

[626] Hossein Haghbin and Mehdi Maadooliat. A journey from univariate to multivariate functional time series: a comprehensive review. WIREs Computational Statistics, 16(2):e1640:1–e1640:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL

https://wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.
1640.

Wang:2024:BDM

[627] Ching-Yun Wang, Wen-Han Hwang, and Xiao Song. Biomarker data with measurement error in medical research: a literature review. WIREs Computational Statistics, 16(2):e1641:1–e1641:??, January– February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1641.

Wunsch:2024:RSM

[628] Milena Wünsch, Christina Sauer, Patrick Callahan, Ludwig Christian Hinske, and Anne-Laure Boulesteix. From RNA sequencing measurements to the final results: a practical guide to navigating the choices and uncertainties of gene set analysis. WIREs Computational Statistics, 16(2): e1643:1–e1643:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1643.

Karakaya:2024:SPQ

[629] Aykut Karakaya and Ahmet Ulu. A survey on post-quantum based approaches for edge computing security. WIREs Computational Statistics, 16(2):e1644:1–e1644:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1644.

Sung:2024:RCM

[630] Chih-Li Sung and Rui Tuo. A review on computer model calibration. WIREs Computational Statistics, 16(2):e1645:1–e1645:??, January– February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1645.

Peng:2024:LSP

[631] Guan-Ju Peng. Learning the sparse prior: Modern approaches. WIREs Computational Statistics, 16(2):e1646:1–e1646:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

URL https://wires.onlinelibrary.wiley.com/doi/epdf/10.1002/
wics.1646.

Betz:2024:DMM

[632] Margaret A. Betz, Rebecca L. Sharples, and Mark Daniel Ward. The data mine model for accessible partnerships in data science. WIREs Computational Statistics, 16(2):e1642:1–e1642:??, January–February 2024. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

//wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1642.

Anonymous:2024:IIb

[633] Anonymous. Issue information. WIREs Computational Statistics, 16(2): e1620:1–e1620:??, March–April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1620.

Rawat:2024:HSD

[634] Anchal Rawat, Santosh Kumar, and Surender Singh Samant. Hate speech detection in social media: Techniques, recent trends, and future challenges. WIREs Computational Statistics, 16(2):e1648:1–e1648:??, March– April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1648.

Shu:2024:DFB

[635] Min Shu and Ruiqiang Song. Detection of financial bubbles using a log-periodic power law singularity (LPPLS) model. WIREs Computational Statistics, 16(2):e1649:1–e1649:??, March–April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1649.

Bhadra:2024:MTC

[636] Anindya Bhadra, Jyotishka Datta, Nicholas G. Polson, Vadim Sokolov, and Jianeng Xu. Merging two cultures: Deep and statistical learning. WIREs Computational Statistics, 16(2):e1647:1–e1647:??, March– April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.1647.

Hung:2024:CRM

[637] Ying-Chao Hung. Correction to “A review of Monte Carlo and quasi-Monte Carlo sampling techniques”. WIREs Computational Statistics, 16 (1):e1650:1–e1650:??, March–April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1650. See [623].

Nolan:2024:MME

[638] John P. Nolan. Modeling multivariate extremes. WIREs Computational Statistics, 16(1):e1652:1–e1652:??, March–April 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1652.

Li:2024:RAM

[639] Zhengpin Li, Yanxi Hou, and Tiandong Wang. Recent advances on mechanisms of network generation: Community, exchangeability, and scale-free properties. WIREs Computational Statistics, 16(1):e1651:1– e1651:??, March–April 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1651.

Anonymous:2024:IIc

[640] Anonymous. Issue information. WIREs Computational Statistics, 16 (3):e1621:1–e1621:??, May–June 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1621.

Lyons:2024:DEI

[641] Emily P. Hendryx Lyons. The discrete empirical interpolation method in class identification and data summarization. WIREs Computational Statistics, 16(3):e1653:1–e1653:??, May–June 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1653.

Anonymous:2024:IId

[642] Anonymous. Issue information. WIREs Computational Statistics, 16(4): e1622:1–e1622:??, July–August 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.1622.

Garcia-Escudero:2024:RCB

[643] Luis A. García-Escudero and Agustín Mayo-Iscar. Robust clustering based on trimming. WIREs Computational Statistics, 16(4):e1658:1– e1658:??, July–August 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1658.

Collins:2024:HAL

[644] David H. Collins, Aparna V. Huzurbazar, and Richard L. Warr. Highly accelerated life testing (HALT): A review from a statistical perspective. WIREs Computational Statistics, 16(4):e70000:1–e70000:??, July– August 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.70000.

Wang:2024:FDA

[645] Shuoyang Wang, Wanyu Zhang, Guanqun Cao, and Yuan Huang. Functional data analysis using deep neural networks. WIREs Computational Statistics, 16(4):e70001:1–e70001:??, July–August 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.70001.

Anonymous:2024:IIe

[646] Anonymous. Issue information. WIREs Computational Statistics, 16 (5):e1623:1–e1623:??, September–October 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1623.

Brown:2024:CRM

[647] Austin Brown and Galin L. Jones. Convergence rates of Metropolis– Hastings algorithms. WIREs Computational Statistics, 16(5):e70002:1– e70002:??, September–October 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70002.

Hubert:2024:RDA

[648] Mia Hubert, Jakob Raymaekers, and Peter J. Rousseeuw. Robust discriminant analysis. WIREs Computational Statistics, 16(5):e70003:1– e70003:??, September–October 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70003.

Anonymous:2024:IIf

[649] Anonymous. Issue information. WIREs Computational Statistics, 16 (6):e1624:1–e1624:??, November–December 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.1624.

Korhonen:2024:RGL

[650] Pekka Korhonen, Klaus Nordhausen, and Sara Taskinen. A review of generalized linear latent variable models and related computational approaches. WIREs Computational Statistics, 16(6):e70005:1–e70005:??, November–December 2024. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.70005.

Todorov:2024:RPE

[651] Valentin Todorov. The R package ecosystem for robust statistics. WIREs Computational Statistics, 16(6):e70007:1–e70007:??, November– December 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.70007.

Anonymous:2025:IIa

[652] Anonymous. Issue information. WIREs Computational Statistics, 17(1): e1654:1–e1654:??, March 2025. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1654.

Tai:2025:UML

[653] Andy Man Yeung Tai, Alireza Kazemi, Jane J. Kim, Jim Schmeckenbecher, Vanessa Kitchin, Janet Suen, Ryan Moro, and R. Michael Krausz. Utilizing machine learning for early intervention and risk management in the opioid overdose crisis. WIREs Computational Statistics, 17(1):e70008:1–e70008:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70008.

Duker:2025:VAM

[654] Marie-Christine Düker, David S. Matteson, Ruey S. Tsay, and Ines Wilms. Vector AutoRegressive moving average models: a review. WIREs Computational Statistics, 17(1):e70009:1–e70009:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.70009.

Fernandez:2025:RRM

[655] Daniel Fernández, Mireia Estopañan, Benjamin Baumer, and Martí Casals. Reporting of regression models for ordinal responses in sports sciences field: a systematic review. WIREs Computational Statistics, 17(1):e70012:1–e70012:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70012.

Thibaudeau:2025:RMM

[656] Yves Thibaudeau. A review of modern multinomial-derived and partitionbased record-linkage methods. WIREs Computational Statistics, 17(1): e70015:1–e70015:??, March 2025. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.70015.

Weiss:2025:MDU

[657] Christian H. Weiß. The mollified (discrete) uniform distribution and its applications. WIREs Computational Statistics, 17(1):e70016:1–e70016:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.70016.

Bhaduri:2025:RTB

[658] Diya Bhaduri, Daniell Toth, and Scott H. Holan. A review of tree-based methods for analyzing survey data. WIREs Computational Statistics, 17(1):e70010:1–e70010:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70010.

Deng:2025:DSR

[659] Lih-Yuan Deng, Nirman Kumar, Henry Horng-Shing Lu, and Ching-Chi Yang. Developments in secure random number generators. WIREs Computational Statistics, 17(1):e70011:1–e70011:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.70011.

Doherty:2025:CAM

[660] Ultán P. Doherty, Rachel M. McLoughlin, and Arthur White. Challenges and adaptations of model-based clustering for flow and mass cytometry. WIREs Computational Statistics, 17(1):e70017:1–e70017:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.70017.

Chen:2025:SED

[661] Tianle Chen, Ying Tian, and Brian Millen. Surrogate endpoints in drug development: a review of statistical and regulatory perspectives and applications. WIREs Computational Statistics, 17(1):e70014:1–e70014:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (elec-

tronic). URL https://wires.onlinelibrary.wiley.com/doi/epdf/
10.1002/wics.70014.

Ghosh:2025:HTM

[662] Indranil Ghosh and Hon Keung Tony Ng. Hidden truncation models: Theory and applications. WIREs Computational Statistics, 17(1): e70019:1–e70019:??, March 2025. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.70019.

Yu:2025:NNR

[663] Dengdeng Yu and Dehan Kong. Nuclear norm regularization. WIREs Computational Statistics, 17(1):e70013:1–e70013:??, March 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://

wires.onlinelibrary.wiley.com/doi/epdf/10.1002/wics.70013.

Anonymous:2025:IIb

[664] Anonymous. Issue information. WIREs Computational Statistics, 17(2): e1655:1–e1655:??, June 2025. CODEN ???? ISSN 1939-0068 (print),

1939-5108 (electronic). URL https://wires.onlinelibrary.wiley.
com/doi/epdf/10.1002/wics.1655.

Goos:2025:ODF

[665] Peter Goos. OMARS designs for factor screening and response surface experimentation in one step: a review. WIREs Computational Statistics, 17(2):e70018:1–e70018:??, June 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.onlinelibrary.

wiley.com/doi/epdf/10.1002/wics.70018.

Park:2025:BCC

[666] Heewon Park and Satoru Miyano. Breaking the complexity of cancer using computational transcriptomic network biology. WIREs Computational Statistics, 17(2):e70020:1–e70020:??, June 2025. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https://wires.

onlinelibrary.wiley.com/doi/epdf/10.1002/wics.70020.

Schneiter:2025:IAC

[667] Kady Schneiter and Olivia Larsen. Instructional applet collections for statistics education. WIREs Computational Statistics, 17(2):e70023:1– e70023:??, June 2025. CODEN ???? ISSN 1939-0068 (print), 1939-

5108 (electronic). URL https://wires.onlinelibrary.wiley.com/
doi/epdf/10.1002/wics.70023.